### v1

# Completeness of sparse and finite local complexity sequences of translates in $L^{p}(\mathbb{R})$

Abstract.
We construct a nonnegative function $g$ and a set $\Lambda=\{\lambda_{1}<\lambda_{2}<\dots\}$ of positive real numbers such that the system $\{g(x-\lambda_{n})\}_{n=1}^{\infty}$ is complete simultaneously in all $L^{p}(\mathbb{R})$ spaces, $1<p<\infty$, and $\Lambda$ satisfies one of the following two additional conditions: either $\Lambda$ is very sparse, namely, the ratios $\lambda_{n+1}/\lambda_{n}$ tend to $1$ arbitrarily slowly; or the differences $\lambda_{n+1}-\lambda_{n}$ attain only two different values, so $\Lambda$ has finite local complexity. Both constructions are, in a sense, extreme: it is well known that $\Lambda$ cannot be Hadamard lacunary and cannot be contained in any arithmetic progression.
Key words and phrases:
Complete systems, translates
2020 Mathematics Subject Classification:
42A10, 42A65, 46E30
N.L. is supported by ISF Grant No. 1044/21. A.T. is supported by the Foundation for the Advancement of Theoretical Physics and Mathematics “BASIS”
1. Introduction
1.1.
The problem of completeness of translates of a single function in $L^{p}(\mathbb{R})$ spaces, goes back to the classical Wiener’s theorems [Wie32], which characterize the functions whose translates span $L^{1}(\mathbb{R})$ or $L^{2}(\mathbb{R})$. If $g\in L^{1}(\mathbb{R})$ then the system of translates
$\{g(x-\lambda)\},\;\lambda\in\mathbb{R},$ (1.1)
is complete in the space $L^{1}(\mathbb{R})$ if and only if the Fourier transform $\widehat{g}$ has no zeros, while if $g\in L^{2}(\mathbb{R})$ then the system (1.1) is complete in $L^{2}(\mathbb{R})$ if and only if $\widehat{g}(t)\neq 0$ a.e.
Beurling [Beu51] proved that if $g\in(L^{p}\cap L^{1})(\mathbb{R})$, $21<p<$ 2, then the translates (1.1) span $L^{p}(\mathbb{R})$ if the zero set of $\widehat{g}$ has Hausdorff dimension less than $2(p-1)/p$. However, this sufficient condition is not necessary. Moreover, the functions $g$ whose translates span $L^{p}(\mathbb{R})$, $21<p<$ 2, cannot be characterized by the zero set of $\widehat{g}$, see [LO11].
1.2.
It is well known that even a discrete sequence of translates may suffice to span the space $L^{p}(\mathbb{R})$. Let us say that a discrete set $\Lambda\subset\mathbb{R}$ is $p$-generating if there exists a function $g\in L^{p}(\mathbb{R})$ (called a “generator”) such that the system of its $\Lambda$-translates
$\{g(x-\lambda)\},\;\lambda\in\Lambda,$ (1.2)
is complete in the space $L^{p}(\mathbb{R})$.
Which sets $\Lambda$ are $p$-generating? Note that if $\Lambda$ is $p$-generating for some $p$, then it is also $q$-generating for every $q>p$, see [OU16, Section 12.6]. It is thus “more difficult” to span the space $L^{p}(\mathbb{R})$ for smaller values of $p$.
For example, if $p>2$ then the set $\Lambda=\mathbb{Z}$ is $p$-generating, i.e. there is a function $g\in L^{p}(\mathbb{R})$ whose integer translates span $L^{p}(\mathbb{R})$. This fact was proved in [AO96]. The result was improved later: it follows from [FOSZ14, Theorem 3.2] that any unbounded set $\Lambda$, no matter how sparse, is $p$-generating for every $p>2$.
To the contrary, the set of integers $\Lambda=\mathbb{Z}$ is not $2$-generating. Indeed, by a simple argument involving the Fourier transform, one can show that a $2$-generating set $\Lambda$ cannot be contained in any arithmetic progression, see [OU16, Section 11.1].
However, it was proved in [Ole97] that any “small perturbation” of the integers,
$\Lambda=\{n+\alpha_{n}:n\in\mathbb{Z}\},\quad 0\neq\alpha_{n}\to 0\quad(|n|\to+\infty),$ (1.3)
is a $2$-generating set. In particular, there exist uniformly discrete $2$-generating sets. We recall that a set $\Lambda$ is uniformly discrete if the distance between any two distinct points of $\Lambda$ is bounded from below by a positive constant.
The case $21<p<$ 2 seems to be more difficult, mainly due to the absence of Plancherel’s theorem in the corresponding $L^{p}(\mathbb{R})$ spaces. Only in [OU18], using methods of complex analysis, it was shown that if the perturbations $\alpha_{n}$ are exponentially small, then the set $\Lambda$ in (1.3) is $p$-generating for every $p>1$. A different approach, based on a result from [Lan64], was developed in [Lev25], which allows one to construct $p$-generating sets, $p>1$, consisting only of perturbations of the positive integers.
On the other hand, for $p=1$ there are no uniformly discrete generating sets. It was proved in [BOU06] that a set $\Lambda$ is $1$-generating if and only if the Beurling–Malliavin density of $\Lambda$ is infinite (for the definition of this density, see e.g. [OU16, Section 4.7]).
1.3.
The results mentioned above provide a complete characterization of the discrete $p$-generating sets for $p=1$ and for $p>2$. To the contrary, the $p$-generating sets for $1<p\leqslant 2$, remain not well understood. In this case, both the size and the arithmetic structure of the set play an important role.
On one hand, a $p$-generating set, $1<p\leqslant 2$, cannot be too sparse. If a positive real sequence $\{\lambda_{n}\}_{n=1}^{\infty}$ satisfies the Hadamard lacunarity condition $\lambda_{n+1}/\lambda_{n}>c>1$, then it is not $2$-generating (and hence not $p$-generating for all $p\leqslant 2$), see [OU16, Section 11.4].
It is known [Ole98], [NO09] that this lacunarity condition is, in a sense, sharp: for any positive sequence $\varepsilon_{n}$ tending to 00, no matter how slowly, there exists a $2$-generating set $\Lambda=\{\lambda_{n}\}_{n=1}^{\infty}$ of positive real numbers satisfying $\lambda_{n+1}/\lambda_{n}>1+\varepsilon_{n}$ for all $n$. This result yields very sparse $2$-generating sets, having any subexponential growth.
Our first theorem extends the aforementioned result to the whole range of exponents $1<p\leqslant 2$, and moreover, we prove that the “generator” $g$ (namely, the function whose $\Lambda$-translates span the space) can be chosen to be a nonnegative function.
Theorem 1.1.
For any positive sequence $\varepsilon_{n}\to 0$ and any $\lambda_{0}>0$, one can find a nonnegative function $g\in\cap_{p>1}L^{p}(\mathbb{R})$ and a positive real sequence $\{\lambda_{n}\}_{n=1}^{\infty}$ satisfying
$\lambda_{n+1}/\lambda_{n}>1+\varepsilon_{n},\quad n=0,1,2,\dots,$ (1.4)
such that the system $\{g(x-\lambda_{n})\}_{n=1}^{\infty}$ is complete in the space $L^{p}(\mathbb{R})$ for every $p>1$.
1.4.
Next, we recall that not only the size, but also the arithmetic structure of the set is important. It was mentioned above that a $2$-generating set (and hence also a $p$-generating set, $1<p\leqslant 2$) cannot be contained in any arithmetic progression.
One can weaken this condition and allow the successive differences $\lambda_{n+1}-\lambda_{n}$ to attain finitely many different (positive) values. A real sequence $\{\lambda_{n}\}_{n=1}^{\infty}$ satisfying this condition is said to have finite local complexity.
Our second result shows that $p$-generating sets of finite local complexity exist for every $p>1$, and moreover, the successive differences $\lambda_{n+1}-\lambda_{n}$ may attain as few as two different values.
Theorem 1.2.
Let $a,b>0$ be linearly independent over the rationals. Then there exist a real sequence $\{\lambda_{n}\}_{n=1}^{\infty}$ with $\lambda_{n+1}-\lambda_{n}\in\{a,b\}$, and a nonnegative Schwartz function $g$, such that the system $\{g(x-\lambda_{n})\}_{n=1}^{\infty}$ is complete in $L^{p}(\mathbb{R})$ for every $p>1$.
The requirement that $a,b$ be linearly independent over the rationals is crucial, for otherwise the points $\{\lambda_{n}\}$ lie in some arithmetic progression and the result fails.
Note that in Theorem 1.2 the function $g$ belongs to the Schwartz class, so it is both smooth and has fast decay. To the contrary, in Theorem 1.1 the function $g$ can be chosen smooth but in general cannot decay fast, see Section 4.5 below.
1.5.
The rest of the paper is organized as follows. In Section 2 we recall some necessary background and fix notation that will be used throughout the paper.
In Section 3 we review and extend the approach from [Lev25], based on Landau’s classical result [Lan64], that allows one to construct uniformly discrete $p$-generating sets for every $p>1$.
In Section 4 we construct a positive sequence $\{\lambda_{n}\}_{n=1}^{\infty}$ which is $p$-generating for every $p>1$, and the ratios $\lambda_{n+1}/\lambda_{n}$ tend to $1$ arbitrarily slowly (Theorem 1.1).
In Section 5 we construct a sequence $\{\lambda_{n}\}_{n=1}^{\infty}$ which is $p$-generating for every $p>1$, and the differences $\lambda_{n+1}-\lambda_{n}$ attain only two different values (Theorem 1.2).
2. Preliminaries
In this section we recall some necessary background and fix notation that will be used throughout the paper.
2.1.
The Schwartz space $\mathcal{S}(\mathbb{R})$ consists of all infinitely smooth functions $\varphi$ on $\mathbb{R}$ such that for each $n,k\geqslant 0$, the seminorm
$\|\varphi\|_{n,k}:=\sup_{x\in\mathbb{R}}(1+|x|)^{n}|\varphi^{(k)}(x)|$ (2.1)
is finite. It is a topological linear space whose topology is induced by the metric
$d(\varphi,\psi):=\sum_{n,k\geqslant 0}2^{-(n+k)}\frac{\|\varphi-\psi\|_{n,k}}{1+\|\varphi-\psi\|_{n,k}}$ (2.2)
which also makes $\mathcal{S}(\mathbb{R})$ a complete, separable metric space.
A tempered distribution on $\mathbb{R}$ is a linear functional on the Schwartz space $\mathcal{S}(\mathbb{R})$ which is continuous with respect to the metric (2.2). We use $\alpha(\varphi)$ to denote the action of a tempered distribution $\alpha$ on a Schwartz function $\varphi$.
We denote by $\operatorname{supp}(\alpha)$ the closed support of a tempered distribution $\alpha$.
If $\varphi$ is a Schwartz function on $\mathbb{R}$ then we define its Fourier transform by
[…]
The Fourier transform of a tempered distribution $\alpha$ is defined by $\widehat{\alpha}(\varphi)=\alpha(\widehat{\varphi})$.
2.2.
Let $A^{p}(\mathbb{T})$, $1\leqslant p<\infty$, denote the Banach space of Schwartz distributions $\alpha$ on the circle $\mathbb{T}=\mathbb{R}/\mathbb{Z}$ whose Fourier coefficients $\{\widehat{\alpha}(n)\}$, $n\in\mathbb{Z}$, belong to $\ell^{p}(\mathbb{Z})$, endowed with the norm $\|\alpha\|_{A^{p}(\mathbb{T})}:=\|\widehat{\alpha}\|_{\ell^{p}(\mathbb{Z})}$. For $p=1$ this is the classical Wiener algebra $A(\mathbb{T})$ of continuous functions with an absolutely convergent Fourier series.
We also use $A^{p}(\mathbb{R})$, $1\leqslant p<\infty$, to denote the Banach space of tempered distributions $\alpha$ on $\mathbb{R}$ whose Fourier transform $\widehat{\alpha}$ is in $L^{p}(\mathbb{R})$, with the norm $\|\alpha\|_{A^{p}(\mathbb{R})}:=\|\widehat{\alpha}\|_{L^{p}(\mathbb{R})}$.
Note that $A^{1}(\mathbb{T})$ and $A^{1}(\mathbb{R})$ are function spaces, continuously embedded in $C(\mathbb{T})$ and $C_{0}(\mathbb{R})$ respectively. Similarly, for $1<p\leqslant 2$ the space $A^{p}$ (on either $\mathbb{T}$ or $\mathbb{R}$) is a function space, continuously embedded in $L^{q}$, $q=p/(p-1)$, by the Hausdorff-Young inequality. On the other hand, $A^{p}$ is not a function space for $p>2$.
2.3.
If $1\leqslant p<q<\infty$ then we have $A^{p}(\mathbb{T})\subset A^{q}(\mathbb{T})$, and moreover, the inequality
$\|\alpha\|_{A^{q}(\mathbb{T})}\leqslant\|\alpha\|_{A^{p}(\mathbb{T})}$ (2.4)
holds for every $\alpha\in A^{p}(\mathbb{T})$. To the contrary, there exists neither inclusion nor norm inequality between different $A^{p}(\mathbb{R})$ spaces.
If $\alpha\in A^{p}(\mathbb{T})$ and $f\in A(\mathbb{T})$ then the product $\alpha\cdot f$ is well defined, and we have
$\|\alpha\cdot f\|_{A^{p}(\mathbb{T})}\leqslant\|\alpha\|_{A^{p}(\mathbb{T})}\cdot\|f\|_{A(\mathbb{T})}.$ (2.5)
2.4.
We introduce an auxiliary norm $\|{{\cdot}}\|_{{*}}$ on the Schwartz space $\mathcal{S}(\mathbb{R})$, defined by
$\|{{u}}\|_{{*}}:=10\cdot\sup_{x\in\mathbb{R}}(1+x^{2})|\widehat{u}(x)|,\quad u\in\mathcal{S}(\mathbb{R}).$ (2.6)
If $u\in\mathcal{S}(\mathbb{R})$ and $f\in A^{p}(\mathbb{T})$, $1\leqslant p<\infty$, then $f$ may be considered also as a $1$-periodic tempered distribution on $\mathbb{R}$, so the product $u\cdot f$ makes sense and is well defined.
Lemma 2.1.
Let $u\in\mathcal{S}(\mathbb{R})$ and $f\in A^{p}(\mathbb{T})$, $1\leqslant p<\infty$. Then
$\|u\cdot f\|_{A^{p}(\mathbb{R})}\leqslant\|{{u}}\|_{{*}}\|f\|_{A^{p}(\mathbb{T})}.$ (2.7)
For a proof see [LT24, Lemma 2.1].
2.5.
[…]
$\|\Delta_{h}\|_{A^{p}(\mathbb{T})}\leqslant h^{(p-1)/p},\quad 1\leqslant p<\infty.$ (2.8)
[…]
$\|\tau_{h}\|_{A^{p}(\mathbb{T})}\leqslant 3h^{(p-1)/p},\quad 1\leqslant p<\infty,$ (2.9)
which follows from (2.8) and the fact that $\tau_{h}(t)=\Delta_{h}(t+h)+\Delta_{h}(t)+\Delta_{h}(t-h)$.
2.6.
By a trigonometric polynomial we mean a finite sum of the form
$P(t)=\sum_{j}a_{j}e^{2\pi i\sigma_{j}t},\quad t\in\mathbb{R},$ (2.10)
where $\{\sigma_{j}\}$ are distinct real numbers, and $\{a_{j}\}$ are complex numbers.
By the spectrum of $P$ we mean the set $\operatorname{spec}(P):=\{\sigma_{j}:a_{j}\neq 0\}$. We observe that if $P$ has integer spectrum, $\operatorname{spec}(P)\subset\mathbb{Z}$, then $P$ is $1$-periodic, that is, $P(t+1)=P(t)$. In this case, $P$ may be considered also as a function on $\mathbb{T}$.
By the degree of $P$ we mean the number $\deg(P):=\min\{r\geqslant 0:\operatorname{spec}(P)\subset[-r,r]\}$.
3. Landau systems and completeness of weighted exponentials
In this section we review and extend the approach from [Lev25], based on Landau’s classical result [Lan64], which allows one to construct uniformly discrete $p$-generating sets $\Lambda$ for every $p>1$. Moreover, by some additional arguments we prove that the “generator” $g$ can be chosen to be a nonnegative Schwartz function.
3.1.
We begin by introducing some terminology.
Definition 3.1.
A set $\Omega\subset\mathbb{R}$ will be called a Landau set if it has the form
$\Omega=\Omega(L,h)=\bigcup_{|l|\leqslant L}[l-\tfrac{1}{2}h,l+\tfrac{1}{2}h]$ (3.1)
[…]
Thus a Landau set consists of a finite number of disjoint closed intervals of length strictly less than one, whose centers lie at integer points.
Definition 3.2.
A real sequence $\{\lambda_{n}\}_{n=1}^{\infty}$ will be called a Landau system if for every $N$ the system $\{e^{2\pi i\lambda_{n}t}\}$, $n>N$, is complete in $L^{2}(\Omega)$ for every Landau set $\Omega\subset\mathbb{R}$.
Landau [Lan64] proved that for any $\varepsilon>0$ there exists a Landau system $\{\lambda_{n}\}_{n=1}^{\infty}$ such that $|\lambda_{n}-n|<\varepsilon$ for all $n$. Using this result we can easily obtain:
Proposition 3.3.
There exists a Landau system $\{\lambda_{n}\}_{n=1}^{\infty}$ satisfying $\lambda_{n}=n+o(1)$.
Proof.
Indeed, let $\{\chi_{k}\}_{k=1}^{\infty}$ be a sequence dense in the Schwartz space, and fix a sequence of Landau sets $\Omega_{k}=\Omega(L_{k},h_{k})$ such that $L_{k}\to+\infty$ and $h_{k}\to 1$. By Landau’s theorem [Lan64], for each $k$ there is a Landau system $\{\lambda^{(k)}_{n}\}_{n=1}^{\infty}$ satisfying $|\lambda^{(k)}_{n}-n|<k^{-1}$ for all $n$. In particular, for every $N$ the system $\{e^{2\pi i\lambda^{(k)}_{n}t}\}$, $n>N$, is complete in $L^{2}(\Omega_{k})$. This allows us to construct by induction a sequence of trigonometric polynomials
$P_{k}(t)=\sum_{N_{k}<n\leqslant N_{k+1}}c_{n}e^{2\pi i\lambda^{(k)}_{n}t}$ (3.2)
such that $\|P_{k}-\chi_{k}\|_{L^{2}(\Omega_{k})}<k^{-1}$. In turn, we next construct a sequence $\{\lambda_{n}\}_{n=1}^{\infty}$ defined by $\lambda_{n}:=\lambda^{(k)}_{n}$ if $n$ is a positive integer belonging to the interval $(N_{k},N_{k+1}]$. It follows that $|\lambda_{n}-n|<k^{-1}$ for $n>N_{k}$, so that $\lambda_{n}=n+o(1)$.
Let us show that $\{\lambda_{n}\}_{n=1}^{\infty}$ is a Landau system. Indeed, let $\Omega=\Omega(L,h)$ be a Landau set. Then for all sufficiently large $k$ we have $\Omega\subset\Omega_{k}$, hence $\|P_{k}-\chi_{k}\|_{L^{2}(\Omega)}\leqslant k^{-1}$. The sequence $\{P_{k}\}$ is therefore dense in the space $L^{2}(\Omega)$. Moreover, for each $k$, the polynomial $P_{k}$ lies in the linear span of the system $\{e^{2\pi i\lambda_{n}t}\}$, $n>N_{k}$. This implies that for every $N$ the system $\{e^{2\pi i\lambda_{n}t}\}$, $n>N$, is complete in the space $L^{2}(\Omega)$. ∎
3.2.
We now come to the main result of the present section.
Definition 3.4.
We denote by $I_{0}(\mathbb{R})$ the closed linear subspace of the Schwartz space $\mathcal{S}(\mathbb{R})$, that consists of all the functions $\varphi\in\mathcal{S}(\mathbb{R})$ satisfying
$\varphi^{(j)}(n+\tfrac{1}{2})=0,\quad n\in\mathbb{Z},\quad j=0,1,2,\dots$ (3.3)
The space $I_{0}(\mathbb{R})$ is a complete, separable metric space with the metric inherited from the Schwartz space.
Theorem 3.5.
Given any Landau system $\{\lambda_{n}\}_{n=1}^{\infty}$ one can find a nonnegative function $u\in I_{0}(\mathbb{R})$, such that for every $N$ the system
$\{u(t)e^{2\pi i\lambda_{n}t}\},\quad n>N,$ (3.4)
is complete in $I_{0}(\mathbb{R})$. Moreover, $u$ can be chosen so that also $\widehat{u}$ is nonnegative.
This is a version of [Lev25, Theorem 4.2] where in addition both $u$ and $\widehat{u}$ are required to be nonnegative. Note that this extra nonnegativity requirement necessitates us to use here a slightly different definition of the space $I_{0}(\mathbb{R})$ than in [Lev25].
3.3.
Next, we turn to the proof of Theorem 3.5.
Let $J_{0}(\mathbb{R})$ be the linear space consisting of all the smooth, compactly supported functions on $\mathbb{R}$ which vanish in a neighborhood of $\mathbb{Z}+\tfrac{1}{2}$. Equivalently, $J_{0}(\mathbb{R})$ consists of all the smooth functions whose support is contained in some Landau set.
The space $J_{0}(\mathbb{R})$ is a dense subspace of $I_{0}(\mathbb{R})$, see [Lev25, Lemma 4.3].
Lemma 3.6.
Let $\chi\in J_{0}(\mathbb{R})$. Then there is a function $\sigma$ with the following properties:
  1. (i)
    $\sigma\in J_{0}(\mathbb{R})$;
  2. (ii)
    $\sigma(t)>0$ for $t\in\operatorname{supp}(\chi)$;
  3. (iii)
    $\sigma$ and $\widehat{\sigma}$ are both nonnegative functions.
Proof.
Let $\rho$ be an even smooth function on $\mathbb{R}$, with $\rho(t)>0$ for $|t|<\tfrac{1}{2}$, and $\rho(t)=0$ for $|t|\geqslant\tfrac{1}{2}$. By replacing $\rho(t)$ with $(\rho\ast\rho)(2t)$ we may assume that $\widehat{\rho}$ is nonnegative.
[…]
$\sigma(t):=\sum_{|l|\leqslant L}\Big{(}1-\frac{|l|}{L+1}\Big{)}\rho\Big{(}\frac{t-l}{h^{\prime}}\Big{)}.$ (3.5)
Then $\sigma$ is a smooth function supported on $\Omega(L,h^{\prime})$, hence (i) holds. Moreover, $\sigma(t)>0$ for $t\in\Omega(L,h)$, so that also condition (ii) is satisfied.
Lastly, it is obvious that $\sigma$ is a nonnegative function. Moreover,
$\widehat{\sigma}(x)=h^{\prime}\cdot\widehat{\rho}(h^{\prime}x)\sum_{|l|\leqslant L}\Big{(}1-\frac{|l|}{L+1}\Big{)}e^{2\pi ilx},$ (3.6)
and we observe that the sum in (3.6) is the classical $L$’th order Fejér kernel, which is nonnegative. Hence (iii) holds as well and the lemma is established. ∎
Proof of Theorem 3.5.
Let $\{\lambda_{n}\}_{n=1}^{\infty}$ be a Landau system, and choose a sequence $\{\chi_{k}\}_{k=1}^{\infty}$ in $J_{0}(\mathbb{R})$ which is dense in the space $I_{0}(\mathbb{R})$. We construct by induction a sequence of functions $u_{k}\in J_{0}(\mathbb{R})$, where $u_{k}$ and $\widehat{u}_{k}$ are both nonnegative, together with an increasing sequence of positive integers $\{N_{k}\}$, and trigonometric polynomials
$P_{k}(t)=\sum_{N_{k}<n<N_{k+1}}c_{n}e^{2\pi i\lambda_{n}t}$ (3.7)
in the following way. We begin by setting $u_{0}:=0$ and $N_{0}:=0$.
At the $k$’th step of the induction, we apply Lemma 3.6 with $\chi=\chi_{k}$ and obtain a function $\sigma_{k}$. Let $u_{k}:=u_{k-1}+\delta_{k}\sigma_{k}$, where $\delta_{k}>0$ is chosen small enough so that
$d(u_{k},u_{k-1})\leqslant 2^{-k},\quad\max_{1\leqslant l\leqslant k-1}d(u_{k}\cdot P_{l},u_{k-1}\cdot P_{l})\leqslant k^{-1}2^{-k}.$ (3.8)
We claim that there exists a polynomial $P_{k}$ as in (3.7), such that $d(u_{k}\cdot P_{k},\chi_{k})<k^{-1}$. Indeed, suppose that $\alpha$ is a tempered distribution on $\mathbb{R}$, which annihilates the system $\{u_{k}(t)e^{2\pi i\lambda_{n}t}\}$, $n>N_{k}$. This means that $\alpha\cdot u_{k}$ satisfies the condition $(\alpha\cdot u_{k})^{\wedge}(-\lambda_{n})=0$ for all $n>N_{k}$. Since $u_{k}\in J_{0}(\mathbb{R})$, the distribution $\alpha\cdot u_{k}$ is supported on some Landau set $\Omega_{k}$. Since $\{\lambda_{n}\}_{n=1}^{\infty}$ is a Landau system, this implies that $\alpha\cdot u_{k}=0$, see [Lev25, Corollary 3.2]. On the other hand, we note that $u_{k}(t)>0$ for $t\in\operatorname{supp}(\chi_{k})$, so we can write $\chi_{k}=u_{k}\cdot v_{k}$ where $v_{k}$ is a smooth function of compact support. Hence
$\alpha(\chi_{k})=\alpha(u_{k}\cdot v_{k})=(\alpha\cdot u_{k})(v_{k})=0.$ (3.9)
By duality, this implies that the function $\chi_{k}$ must lie in the closed linear subspace of $I_{0}(\mathbb{R})$ spanned by the system $\{u_{k}(t)e^{2\pi i\lambda_{n}t}\}$, $n>N_{k}$. As a consequence, we conclude that indeed there is a polynomial (3.7) satisfying $d(u_{k}\cdot P_{k},\chi_{k})<k^{-1}$.
The sequence $\{u_{k}\}$ converges in the space $I_{0}(\mathbb{R})$ to some function $u\in I_{0}(\mathbb{R})$ such that both $u$ and $\widehat{u}$ are nonnegative. We have
$d(u\cdot P_{k},\chi_{k})\leqslant d(u_{k}\cdot P_{k},\chi_{k})+\sum_{l=k+1}^{\infty}d(u_{l}\cdot P_{k},u_{l-1}\cdot P_{k})<2k^{-1}$ (3.10)
for every $k$, due to (3.8). As a consequence, the sequence $\{u\cdot P_{k}\}$ is dense in $I_{0}(\mathbb{R})$. Moreover, $u\cdot P_{k}$ belongs to the linear span of the system $\{u(t)e^{2\pi i\lambda_{n}t}\}$, $n>N_{k}$, due to (3.7). This implies that for every $N$ the system (3.4) is complete in $I_{0}(\mathbb{R})$. ∎
3.4.
As a consequence of the result just proved, we obtain:
Corollary 3.7.
Given any Landau system $\{\lambda_{n}\}_{n=1}^{\infty}$ one can find a nonnegative Schwartz function $g$ on $\mathbb{R}$, such that for every $N$ the system
$\{g(x-\lambda_{n})\},\quad n>N,$ (3.11)
is complete in the space $L^{p}(\mathbb{R})$ for every $p>1$.
Indeed, this follows from Theorem 3.5 and the fact that the Fourier transform maps $I_{0}(\mathbb{R})$ continuously and densely into $L^{p}(\mathbb{R})$, $p>1$, see [Lev25, Section 5].
In particular, by applying Corollary 3.7 to a Landau system $\{\lambda_{n}\}_{n=1}^{\infty}$ satisfying the condition $\lambda_{n}=n+o(1)$, we obtain [Lev25, Theorem 1.1] with the extra property that the “generator” $g$ is taken to be nonnegative.
4. Sparse complete sequences of translates
4.1.
In this section we prove Theorem 1.1. First, by using the fact that the Fourier transform is an isometric isomorphism $A^{p}(\mathbb{R})\to L^{p}(\mathbb{R})$, we can reformulate Theorem 1.1 as a result about completeness of weighted exponentials in $A^{p}(\mathbb{R})$.
Theorem 4.1.
For any positive sequence $\varepsilon_{n}\to 0$ and any $\lambda_{0}>0$, there is a nonnegative $w\in\cap_{p>1}A^{p}(\mathbb{R})$ with $\widehat{w}$ nonnegative, and a positive real sequence $\{\lambda_{n}\}_{n=1}^{\infty}$ satisfying
$\lambda_{n+1}/\lambda_{n}>1+\varepsilon_{n},\quad n=0,1,2,\dots,$ (4.1)
such that the system $\{w(t)e^{2\pi i\lambda_{n}t}\}_{n=1}^{\infty}$ is complete in the space $A^{p}(\mathbb{R})$ for every $p>1$.
Theorem 1.1 is obtain as a consequence of this result, by taking $g=\widehat{w}$.
The proof of Theorem 4.1 given below combines the approach in [NO09] together with techniques from [Lev25] and [LT24].
4.2.
Lemma 4.2.
Given any $60<h<1/$ 6, one can find a nonnegative function $\varphi\in A(\mathbb{T})$ with the following properties:
  1. (i)
    $\widehat{\varphi}(0)=1$, $\widehat{\varphi}(n)\geqslant 0$ for all $n\in\mathbb{Z}$;
  2. (ii)
[…]
  3. (iii)
    $\|\varphi-1\|_{A^{p}(\mathbb{T})}\leqslant 6\cdot h^{(p-1)/p}$ for every $p\geqslant 1$.
Proof.
It can be verified in a similar way to [LT24, Lemma 5.3] that the function
$\varphi(t):=1+3\cdot\Delta_{h}(t)-\tau_{h}(t+1/2)$ (4.2)
satisfies the required properties. ∎
Lemma 4.3.
Given $p>1$ and $\varepsilon>0$ there exist two trigonometric polynomials $P$ and $\gamma$ with integer spectrum, such that
  1. (i)
    $\gamma(t)>0$ for all $t\in\mathbb{T}$ (in particular, $\gamma$ has no zeros);
  2. (ii)
    $\widehat{\gamma}(0)=1$, $\widehat{\gamma}(n)\geqslant 0$ for all $n\in\mathbb{Z}$;
  3. (iii)
    $\|\gamma-1\|_{A^{p}(\mathbb{T})}<\varepsilon$;
  4. (iv)
    $\widehat{P}(-n)=0$ for $n=0,1,2,\dots$;
  5. (v)
    $\|P\cdot\gamma-1\|_{A^{p}(\mathbb{T})}<\varepsilon$.
Proof.
We choose and fix a small $h=h(p,\varepsilon)>0$. Let $\varphi$ be the function given by Lemma 4.2, and let $\chi(t):=1-\tau_{2h}(t+1/2)$. We first claim that one can find a trigonometric polynomial $P$ with integer spectrum satisfying the condition (iv) and such that $\|P\cdot\varphi-\chi\|_{A^{p}(\mathbb{T})}<\varepsilon/2$.
Indeed, let $\alpha$ be a Schwartz distribution belonging to the dual space $(A^{p}(\mathbb{T}))^{*}=A^{q}(\mathbb{T})$, $q=p/(p-1)$, and suppose that $\alpha$ annihilates the system
$\{\varphi(t)e^{2\pi int}\},\quad n=1,2,3,\dots$ (4.3)
[…]
$\alpha(\chi)=\alpha(\varphi\cdot\psi\cdot\chi)=(\alpha\cdot\varphi)(\psi\cdot\chi)=0.$ (4.4)
By duality, this implies that $\chi$ must lie in the closed linear subspace of $A^{p}(\mathbb{T})$ spanned by the system (4.3), so there is a trigonometric polynomial $P$ with integer spectrum satisfying (iv) and such that $\|P\cdot\varphi-\chi\|_{A^{p}(\mathbb{T})}<\varepsilon/2$.
If $h=h(p,\varepsilon)>0$ is small enough, then both $\|\varphi-1\|_{A^{p}(\mathbb{T})}$ and $\|\chi-1\|_{A^{p}(\mathbb{T})}$ do not exceed $\varepsilon/2$. As a consequence, $\|P\cdot\varphi-1\|_{A^{p}(\mathbb{T})}<\varepsilon$. We can thus conclude the proof by choosing $\gamma$ to be a Fejér sum of $\varphi$ of sufficiently high order, that is, $\gamma=\varphi\ast K_{N}$ where $K_{N}$ is the classical $N$’th order Fejér kernel and $N$ is sufficiently large. ∎
4.3.
Lemma 4.4.
Let $u\in\mathcal{S}(\mathbb{R})$, let $H$ be a trigonometric polynomial (with real spectrum), and let $p>1$ and $\eta>0$ be given. Then there is $\delta>0$ such that for any $d>0$ one can find trigonometric polynomials $G$ and $Q$ satisfying the following properties:
  1. (i)
    $\operatorname{spec}(G)\subset\mathbb{Z}$;
  2. (ii)
    $G(t)>0$ for all $t\in\mathbb{T}$ (in particular, $G$ has no zeros);
  3. (iii)
    $\widehat{G}(0)=1$, $\widehat{G}(n)\geqslant 0$ for all $n\in\mathbb{Z}$;
  4. (iv)
    $\|G-1\|_{A^{p}(\mathbb{T})}<\eta$;
  5. (v)
    $\operatorname{spec}(Q)$ is contained in some set $\{\lambda_{1},\lambda_{2},\dots,\lambda_{N}\}\subset\mathbb{R}$;
  6. (vi)
    $\lambda_{1}>d$, and $\lambda_{j+1}/\lambda_{j}>1+\delta$ for $j=1,2,\dots,N-1$;
  7. (vii)
    $\|u\cdot(G\cdot Q-H)\|_{A^{q}(\mathbb{R})}<\eta$ for every $q\geqslant p$.
Proof.
Let us denote $H(t)=\sum_{n=1}^{K}c_{n}e^{2\pi i\sigma_{n}t}$, where $\{\sigma_{n}\}$ are distinct real numbers, and $\{c_{n}\}$ are complex numbers. We choose a small $\varepsilon=\varepsilon(u,H,p,\eta)>0$, and apply Lemma 4.3 to obtain trigonometric polynomials $P$ and $\gamma$. Next, we choose large positive integers $\nu_{1}<\nu_{2}<\dots<\nu_{K}$ and set
$G(t)=\prod_{n=1}^{K}\gamma(\nu_{n}t),\quad Q(t)=\sum_{n=1}^{K}c_{n}e^{2\pi i\sigma_{n}t}P(\nu_{n}t).$ (4.5)
It is obvious that the properties (i) and (ii) are satisfied.
If $\nu_{1},\dots,\nu_{K}$ are chosen sufficiently fast increasing, then also (iii) holds. Moreover,
$\|G-1\|_{A^{p}(\mathbb{T})}^{p}=\|G\|_{A^{p}(\mathbb{T})}^{p}-1=\|\gamma\|_{A^{p}(\mathbb{T})}^{pK}-1<(1+\varepsilon)^{pK}-1<\eta^{p}$ (4.6)
provided that $\varepsilon$ is chosen small enough, so we obtain (iv).
Let $M=M(u,H)=1+\sum_{n=1}^{K}\|{{c_{n}u(t)e^{2\pi i\sigma_{n}t}}}\|_{{*}}$. We claim that
$\|G(t)P(\nu_{n}t)-1\|_{A^{p}(\mathbb{T})}<M^{-1}\eta,\quad n=1,2,\dots,K,$ (4.7)
for $\varepsilon$ small enough. Indeed, we have
$G(t)P(\nu_{n}t)-1=G_{n}(t)(\gamma(\nu_{n}t)P(\nu_{n}t)-1)+(G_{n}(t)-1),$ (4.8)
where $G_{n}(t):=\prod_{j\neq n}\gamma(\nu_{j}t)$. The norm of the last summand in (4.8) can be estimated similarly to (4.6), so we can assume that $\|G_{n}-1\|_{A^{p}(\mathbb{T})}<(2M)^{-1}\eta$ by choosing $\varepsilon$ small enough. If $\nu_{1},\dots,\nu_{K}$ are chosen sufficiently fast increasing, then the norm of the first summand on the right hand side of (4.8) is equal to
$\|G_{n}\|_{A^{p}(\mathbb{T})}\|\gamma\cdot P-1\|_{A^{p}(\mathbb{T})}\leqslant 2\varepsilon<(2M)^{-1}\eta$ (4.9)
for $\varepsilon$ small enough, and thus (4.7) follows.
Next we verify property (vii). Indeed, we have
$u(t)(G(t)Q(t)-H(t))=\sum_{n=1}^{K}c_{n}u(t)e^{2\pi i\sigma_{n}t}(G(t)P(\nu_{n}t)-1),$ (4.10)
hence using (2.4), (2.7), (4.7) we obtain for every $q\geqslant p$,
$\|u\cdot(G\cdot Q-H)\|_{A^{q}(\mathbb{R})}\leqslant\sum_{n=1}^{K}\|{{c_{n}u(t)e^{2\pi i\sigma_{n}t}}}\|_{{*}}\cdot\|G(t)P(\nu_{n}t)-1\|_{A^{p}(\mathbb{T})}<\eta,$ (4.11)
so condition (vii) is satisfied.
Lastly we turn to establish the properties (v) and (vi). We denote $L=\deg(P)$, and take $\delta=(1+L)^{-1}$. Now suppose that $d>0$ is given. If $\nu_{1},\dots,\nu_{K}$ are chosen sufficiently fast increasing, then we have
$\operatorname{spec}(Q)\subset\bigcup_{n=1}^{K}J_{n},\quad J_{n}=\sigma_{n}+\{\nu_{n},2\nu_{n},3\nu_{n},\dots,L\nu_{n}\},$ (4.12)
and $J_{n+1}$ follows $J_{n}$ for each $n=1,2,\dots,K-1$. Let us write
$\bigcup_{n=1}^{K}J_{n}=\{\lambda_{1},\lambda_{2},\dots,\lambda_{N}\},$ (4.13)
where $\lambda_{j}$ are ordered increasingly. Then $\lambda_{1}=\sigma_{1}+\nu_{1}>d$, provided that $\nu_{1}$ is large enough (we note that $\nu_{1},\dots,\nu_{K}$ are allowed to depend on $d$). Next, suppose that $\lambda_{j}$ and $\lambda_{j+1}$ are two consecutive elements of the set (4.13). If $\lambda_{j}$ and $\lambda_{j+1}$ lie in the same block $J_{n}$, then according to (4.12) we have
$\frac{\lambda_{j+1}}{\lambda_{j}}=\frac{\sigma_{n}+(l+1)\nu_{n}}{\sigma_{n}+l\nu_{n}}$ (4.14)
for some $l\in\{1,2,\dots,L-1\}$. Otherwise, $\lambda_{j}$ is the last element of $J_{n}$, while $\lambda_{j+1}$ is the first element of $J_{n+1}$, for some $n\in\{1,2,\dots,K-1\}$. In this case,
$\frac{\lambda_{j+1}}{\lambda_{j}}=\frac{\sigma_{n+1}+\nu_{n+1}}{\sigma_{n}+L\nu_{n}}.$ (4.15)
In either case, (4.14) or (4.15), we can ensure by taking $\nu_{1},\dots,\nu_{K}$ large enough that the condition $\lambda_{j+1}/\lambda_{j}>1+\delta$ holds for $j=1,2,\dots,N-1$. Thus properties (v) and (vi) are satisfied and the lemma is proved. ∎
4.4. Proof of Theorem 4.1
The proof consists of several steps.
4.4.1.
Let $0<\varepsilon_{n}\to 0$ be given, and we may also assume that $\varepsilon_{n+1}<\varepsilon_{n}$ for each $n$. We start by choosing a sequence $\{\chi_{k}\}_{k=1}^{\infty}\subset I_{0}(\mathbb{R})$ which is dense in the space $I_{0}(\mathbb{R})$.
By Proposition 3.3 there exists a Landau system $\{\sigma_{n}\}_{n=1}^{\infty}$ satisfying
$\sigma_{n}=n+o(1),\quad n\to+\infty.$ (4.16)
By an application of Theorem 3.5 we obtain a nonnegative function $u_{0}\in I_{0}(\mathbb{R})$ with $\widehat{u}_{0}$ nonnegative, such that for every $N$ the system $\{u_{0}(t)e^{2\pi i\sigma_{n}t}\}$, $n>N$, is complete in $I_{0}(\mathbb{R})$. We will construct by induction a sequence of functions $\{u_{k}\}_{k=1}^{\infty}\subset I_{0}(\mathbb{R})$ such that for each $k$ and every $N$, the system
$\{u_{k}(t)e^{2\pi i\sigma_{n}t}\},\;n>N,$ (4.17)
is complete in the space $I_{0}(\mathbb{R})$.
The sequence $\{\lambda_{n}\}_{n=1}^{\infty}$ satisfying (4.1) will be constructed by the same induction.
At the $k$’th step of the induction, suppose that the elements $\{\lambda_{n}\}$, $1\leqslant n\leqslant N_{k}$, satisfying (4.1) have already been defined. Given any positive integer $l_{k}$ we use the completeness of the system $\{u_{k-1}(t)e^{2\pi i\sigma_{n}t}\}$, $n>l_{k}$, in $I_{0}(\mathbb{R})$ to find a polynomial
$H_{k}(t)=\sum_{l_{k}<n<l^{\prime}_{k}}c_{n,k}e^{2\pi i\sigma_{n}t}$ (4.18)
such that
$d(u_{k-1}\cdot H_{k},\chi_{k})<k^{-1}.$ (4.19)
We choose a small number $\eta_{k}>0$ so that
$\eta_{k}\cdot\Big{(}1+\|{{u_{k-1}}}\|_{{*}}+\sum_{j=1}^{k-1}\|{{u_{k-1}\cdot Q_{j}}}\|_{{*}}\Big{)}<2^{-k}k^{-1},$ (4.20)
and let $p_{k}:=1+k^{-1}$. Now invoke Lemma 4.4 with $u_{k-1}$, $H_{k}$, $p_{k}$ and $\eta_{k}$ to obtain a number $\delta_{k}>0$. We choose $M_{k}$ large enough so that we have $\varepsilon_{M_{k}}<\delta_{k}$, and add arbitrary elements $\{\lambda_{n}\}$, $N_{k}<n\leqslant M_{k}$, while keeping the condition (4.1). Now set $d_{k}:=(1+\varepsilon_{M_{k}})\lambda_{M_{k}}$ and use Lemma 4.4 to obtain trigonometric polynomials $G_{k}$ and $Q_{k}$. The spectrum of $Q_{k}$ is contained in a sequence $\{\lambda_{n}\}$, $M_{k}<n\leqslant N_{k+1}$, which still satisfies (4.1) thanks to Lemma 4.4(v),(vi).
Finally, define $u_{k}:=u_{k-1}\cdot G_{k}$ which is a nonnegative function in $I_{0}(\mathbb{R})$. Recall that by the previous inductive step, for every $N$ the system $\{u_{k-1}(t)e^{2\pi i\sigma_{n}t}\}$, $n>N$, is complete in $I_{0}(\mathbb{R})$. Since multiplication by a trigonometric polynomial $G_{k}$ with no zeros maps $I_{0}(\mathbb{R})$ continuously and densely into $I_{0}(\mathbb{R})$, it follows that also the system (4.17) is complete in the space $I_{0}(\mathbb{R})$. This concludes the inductive construction.
4.4.2.
Since we have $1<p_{k}\to 1$, then for each $p>1$ there exists a sufficiently large $k(p)$ such that $1<p_{k}<p$ for all $k\geqslant k(p)$. For such $k$, using (2.4), (2.7) we have
$\|u_{k}-u_{k-1}\|_{A^{p}(\mathbb{R})}=\|u_{k-1}(G_{k}-1)\|_{A^{p}(\mathbb{R})}\leqslant\|{{u_{k-1}}}\|_{{*}}\|G_{k}-1\|_{A^{p_{k}}(\mathbb{T})}.$ (4.21)
Using Lemma 4.4(iv) and (4.20) we conclude that
$\|u_{k}-u_{k-1}\|_{A^{p}(\mathbb{R})}<2^{-k},\quad k\geqslant k(p),$ (4.22)
which implies that the sequence $u_{k}$ converges in the space $A^{p}(\mathbb{R})$ for every $p>1$.
The limit of the sequence $u_{k}$ is thus a nonnegative function $w\in\cap_{p>1}A^{p}(\mathbb{R})$. We recall that $u_{0}$ has a nonnegative Fourier transform $\widehat{u}_{0}$, while each $G_{k}$ is a trigonometric polynomial with nonnegative Fourier coefficients. This implies that $\widehat{u}_{k}$ is nonnegative for each $k$, and as a consequence, $\widehat{w}$ is a nonnegative function in $\cap_{p>1}L^{p}(\mathbb{R})$.
4.4.3.
Next, we fix $p>1$ and claim that for every $N$ the system
$\{w(t)e^{2\pi i\lambda_{n}t}\},\quad n>N,$ (4.23)
is complete in the space $A^{p}(\mathbb{R})$. To prove this, it will be enough to show that
$\|w\cdot Q_{k}-\chi_{k}\|_{A^{p}(\mathbb{R})}\to 0,\quad k\to+\infty.$ (4.24)
Indeed, the sequence $\{\chi_{k}\}_{k=1}^{\infty}$ is dense in the space $I_{0}(\mathbb{R})$, while $I_{0}(\mathbb{R})$ is continuously and densely embedded in $A^{p}(\mathbb{R})$, see [Lev25, Section 5]. Hence $\{\chi_{k}\}_{k=1}^{\infty}$ is dense also in the space $A^{p}(\mathbb{R})$. Recalling that $\operatorname{spec}(Q_{k})$ is contained in the sequence $\{\lambda_{n}\}$, $n>N_{k}$, the condition (4.24) implies the completeness of the system (4.23) in $A^{p}(\mathbb{R})$.
In order to establish (4.24) we write
$w\cdot Q_{k}-\chi_{k}=(u_{k-1}\cdot H_{k}-\chi_{k})+u_{k-1}\cdot(G_{k}\cdot Q_{k}-H_{k})+(w-u_{k})\cdot Q_{k},$ (4.25)
and estimate the $A^{p}(\mathbb{R})$ norm of each term on the right hand side of (4.25).
To estimate the norm of the first term, we recall that the space $I_{0}(\mathbb{R})$ is continuously embedded in $A^{p}(\mathbb{R})$, so (4.19) implies that
$\|u_{k-1}\cdot H_{k}-\chi_{k}\|_{A^{p}(\mathbb{R})}\to 0,\quad k\to+\infty.$ (4.26)
To estimate the norm of the second term we recall that by Lemma 4.4(vii) we have
$\|u_{k-1}\cdot(G_{k}\cdot Q_{k}-H_{k})\|_{A^{p}(\mathbb{R})}<\eta_{k},\quad k\geqslant k(p).$ (4.27)
It remains to estimate the last term in (4.25). We note that for any fixed $k$ we have $u_{j}\cdot Q_{k}\to w\cdot Q_{k}$ as $j\to+\infty$ in the $A^{p}(\mathbb{R})$ norm, hence for $k\geqslant k(p)$ we have
$\|(w-u_{k})\cdot Q_{k}\|_{A^{p}(\mathbb{R})}\leqslant\sum_{j=k+1}^{\infty}\|(u_{j}-u_{j-1})\cdot Q_{k}\|_{A^{p}(\mathbb{R})}$ (4.28)
$\qquad=\sum_{j=k+1}^{\infty}\|u_{j-1}\cdot(G_{j}-1)\cdot Q_{k}\|_{A^{p}(\mathbb{R})}$ (4.29)
$\qquad\leqslant\sum_{j=k+1}^{\infty}\|{{u_{j-1}\cdot Q_{k}}}\|_{{*}}\cdot\|G_{j}-1\|_{A^{p_{j}}(\mathbb{T})}$ (4.30)
$\qquad\leqslant\sum_{j=k+1}^{\infty}2^{-j}j^{-1}<k^{-1},$ (4.31)
where the inequality (4.30) follows from (2.4), (2.7), while to obtain the inequality (4.31) we have used Lemma 4.4(iv) and (4.20). Finally, combining (4.25), (4.26), (4.27) and (4.28)–(4.31) yields the required estimate (4.24), and so we establish the completeness of the system (4.23) in $A^{p}(\mathbb{R})$ for every $p>1$.
This completes the proof of Theorem 4.1, and as a consequence, Theorem 1.1 is also established. ∎
4.5. Remarks
[…]
$M_{n}:=\sup_{x\in\mathbb{R}}\sum_{j\in\mathbb{Z}}(1+|x-j|^{n})u_{0}(x-j)$ (4.32)
for each $k$, and the desired conclusion follows by passing to the limit as $k\to+\infty$.
2. On the other hand, if $\{\varepsilon_{n}\}$ decrease slowly, then the function $g$ cannot have fast decay, in fact, $g$ cannot be chosen in $L^{1}(\mathbb{R})$. For otherwise, $\widehat{g}$ would be a continuous function and the exponential system $\{e^{2\pi i\lambda_{n}t}\}$ would be complete on some interval of positive length, which is not possible if $\{\lambda_{n}\}$ is sparse, see e.g. [OU16, Section 4.7].
3. It can be inferred from the proof of Lemma 4.4 that by choosing $l_{k}\to+\infty$ and using (4.16), (4.18), we can obtain a sequence $\{\lambda_{n}\}$ such that $\operatorname{dist}(\lambda_{n},\mathbb{Z})\to 0$.
5. Finite local complexity sequences of translates
5.1.
In this section we prove Theorem 1.2. The main ingredients of the proof consist of Theorem 3.5 (or its consequence Corollary 3.7), and the following lemma.
Lemma 5.1.
Let $a$ be an irrational positive real number, $\Omega=\Omega(L,h)$ be a Landau set given by (3.1), $\chi$ be a continuous function on $\Omega$, and let $\lambda_{0}\in\mathbb{R}$ and $\varepsilon>0$ be given. Then there exists a trigonometric polynomial $P(t)=\sum_{n=1}^{K}c_{n}e^{2\pi i\lambda_{n}t}$ such that
  1. (i)
    $|P(t)-\chi(t)|\leqslant\varepsilon$ for all $t\in\Omega$;
  2. (ii)
    $\lambda_{n+1}-\lambda_{n}\in\{1,a\}$ for $n=0,1,\dots,K-1$.
Proof.
By multiplying $\chi(t)$ on the exponential $e^{-2\pi i\lambda_{0}t}$ we may assume that $\lambda_{0}=0$. We will construct a trigonometric polynomial $P$ satisfying (i) which is of the form
$P(t)=\sum_{k=0}^{2L}e^{2\pi ikat}Q_{k}(t),$ (5.1)
where $Q_{k}$ are trigonometric polynomials with integer spectrum such that
$\operatorname{spec}(Q_{k})\subset\{j:N_{k}\leqslant j\leqslant N_{k+1}\},\quad 0\leqslant k\leqslant 2L,$ (5.2)
and $\{N_{k}\}$, $10\leqslant k\leqslant 2L+$ 1, is an increasing sequence of integers with $N_{0}=1$. Then
$\operatorname{spec}(P)\subset\bigcup_{k=0}^{2L}J_{k},\quad J_{k}=\{ka+j:\;N_{k}\leqslant j\leqslant N_{k+1}\},$ (5.3)
and $J_{k+1}$ follows $J_{k}$ for each $k=0,1,\dots,2L-1$. Let us write
$\bigcup_{k=0}^{2L}J_{k}=\{\lambda_{1},\lambda_{2},\dots,\lambda_{K}\},$ (5.4)
where $\lambda_{n}$ are ordered increasingly. Then $\lambda_{1}-\lambda_{0}=1$. Next, suppose that $\lambda_{n}$ and $\lambda_{n+1}$ are two consecutive elements of the set (5.4). If $\lambda_{n}$ and $\lambda_{n+1}$ lie in the same block $J_{k}$, then $\lambda_{n+1}-\lambda_{n}=1$. Otherwise, $\lambda_{n}$ is the last element of $J_{k}$, while $\lambda_{n+1}$ is the first element of $J_{k+1}$, for some $k\in\{0,1,\dots,2L-1\}$, and in this case $\lambda_{n+1}-\lambda_{n}=a$. We thus see that condition (ii) is satisfied.
We now turn to the construction of the polynomial $P$. First, we note that (5.1) and the periodicity of $Q_{k}$ imply that for any integer $l$ we have
$P(t+l)=\sum_{k=0}^{2L}e^{2\pi ikla}H_{k}(t),\quad H_{k}(t):=e^{2\pi ikat}Q_{k}(t).$ (5.5)
If we write (5.5) for $|l|\leqslant L$, then we obtain a system of $12L+$ 1 linear equations for the polynomials $H_{k}$, $0\leqslant k\leqslant 2L$, with a Vandermonde coefficient matrix $\{e^{2\pi ikla}\}$ whose determinant does not vanish due to the irrationality of $a$. By solving this system we get
$H_{k}(t)=\sum_{|l|\leqslant L}d_{kl}P(t+l),\quad 0\leqslant k\leqslant 2L,$ (5.6)
where $\{d_{kl}\}$ are the entries of the inverse matrix, which depend on $a$ and $L$ only.
Now recall that $10<h<$ 1, and hence for every $N$ the exponential system $\{e^{2\pi int}\}$, $n>N$, is complete in the space $C[-\tfrac{1}{2}h,\tfrac{1}{2}h]$, see e.g. [You01, Section 3.1]. Using this fact, we can choose the trigonometric polynomials $Q_{0},Q_{1},\ldots,Q_{2L}$ one after another, so that their spectra “follow one another”, i.e.  satisfy (5.2), and such that
$\max_{|t|\leqslant\frac{1}{2}h}\Big{|}Q_{k}(t)-e^{-2\pi ikat}\sum_{|l|\leqslant L}d_{kl}\chi(t+l)\Big{|}\leqslant\delta,\quad 0\leqslant k\leqslant 2L,$ (5.7)
where $\delta:=\varepsilon\cdot(2L+1)^{-1}$. Having chosen the polynomials $Q_{k}$, we next define the polynomial $P$ using (5.1). It now follows from (5.5), (5.6), (5.7) that if we denote
$E_{k}(t):=\sum_{|l|\leqslant L}d_{kl}(P(t+l)-\chi(t+l)),\quad t\in[-\tfrac{1}{2}h,\tfrac{1}{2}h],\quad 0\leqslant k\leqslant 2L,$ (5.8)
then $|E_{k}(t)|\leqslant\delta$. In turn, this implies that for $|l|\leqslant L$ and $t\in[-\tfrac{1}{2}h,\tfrac{1}{2}h]$ we have
$|P(t+l)-\chi(t+l)|=\Big{|}\sum_{k=0}^{2L}e^{2\pi ikla}E_{k}(t)\Big{|}\leqslant(2L+1)\cdot\delta=\varepsilon,$ (5.9)
that is, $|P(t)-\chi(t)|\leqslant\varepsilon$ for all $t\in\Omega$. Thus the lemma is proved. ∎
5.2. Proof of Theorem 1.2
By rescaling, it would be enough to prove the result in the case where one of the numbers $a,b$ is equal to $1$. Hence, in what follows we shall assume that $b=1$, and $a$ is an irrational positive number (since $a,b$ are linearly independent over the rationals).
By virtue of Corollary 3.7, it would suffice that we prove the existence of a Landau system $\{\lambda_{n}\}_{n=1}^{\infty}$ satisfying $\lambda_{n+1}-\lambda_{n}\in\{1,a\}$ for all $n$. Let $\{\chi_{k}\}_{k=1}^{\infty}$ be a sequence which is dense in the Schwartz space, and fix a sequence of Landau sets $\Omega_{k}=\Omega(L_{k},h_{k})$ such that $L_{k}\to+\infty$ and $h_{k}\to 1$. We now construct by induction an increasing sequence of positive integers $\{N_{k}\}$, and trigonometric polynomials
$P_{k}(t)=\sum_{N_{k}<n\leqslant N_{k+1}}c_{n}e^{2\pi i\lambda_{n}t}$ (5.10)
in the following way. At the $k$’th step of the induction, we apply Lemma 5.1 with $\Omega_{k}$, $\chi_{k}$ and $\varepsilon_{k}=k^{-1}$, in order to obtain a trigonometric polynomial $P_{k}$ of the form (5.10) such that $|P_{k}(t)-\chi_{k}(t)|\leqslant k^{-1}$ for all $t\in\Omega_{k}$. Moreover, Lemma 5.1 allows us to choose the frequencies $\{\lambda_{n}\}_{n=1}^{\infty}$ so that $\lambda_{n+1}-\lambda_{n}\in\{1,a\}$ for all $n$.
It remains to show that $\{\lambda_{n}\}_{n=1}^{\infty}$ is a Landau system. This can be done in the same way as in the proof of Proposition 3.3. Indeed, let $\Omega=\Omega(L,h)$ be a Landau set. Then for all sufficiently large $k$ we have $\Omega\subset\Omega_{k}$, hence $|P_{k}(t)-\chi_{k}(t)|\leqslant k^{-1}$ for all $t\in\Omega$. The sequence $\{P_{k}\}$ is therefore dense in the space $L^{2}(\Omega)$. Moreover, for each $k$, the polynomial $P_{k}$ lies in the linear span of the system $\{e^{2\pi i\lambda_{n}t}\}$, $n>N_{k}$. This implies that for every $N$ the system $\{e^{2\pi i\lambda_{n}t}\}$, $n>N$, is complete in the space $L^{2}(\Omega)$. Hence the sequence $\{\lambda_{n}\}_{n=1}^{\infty}$ is a Landau system, and Theorem 1.2 is thus proved. ∎