### v2

# Completeness of sparse, almost integer and finite local complexity sequences of translates in $L^{p}(\mathbb{R})$

Abstract.
A real sequence $\Lambda=\{\lambda_{n}\}_{n=1}^{\infty}$ is called $p$-generating if there exists a function $g$ whose translates $\{g(x-\lambda_{n})\}_{n=1}^{\infty}$ span the space $L^{p}(\mathbb{R})$. While the $p$-generating sets were completely characterized for $p=1$ and $p>2$, the case $1<p\leqslant 2$ remains not well understood. In this case, both the size and the arithmetic structure of the set play an important role. In the present paper, (i) We show that a $p$-generating set $\Lambda$ of positive real numbers can be very sparse, namely, the ratios $\lambda_{n+1}/\lambda_{n}$ may tend to $1$ arbitrarily slowly; (ii) We prove that every “almost integer” sequence $\Lambda$, i.e. satisfying $\lambda_{n}=n+\alpha_{n}$, $0\neq\alpha_{n}\to 0$, is $p$-generating; and (iii) We construct $p$-generating sets $\Lambda$ such that the successive differences $\lambda_{n+1}-\lambda_{n}$ attain only two different positive values. The constructions are, in a sense, extreme: it is well known that $\Lambda$ cannot be Hadamard lacunary and cannot be contained in any arithmetic progression.
Key words and phrases:
Complete systems, translates
2020 Mathematics Subject Classification:
42A10, 42A65, 46E30
N.L. is supported by ISF Grant No. 1044/21. A.T. is supported by the Foundation for the Advancement of Theoretical Physics and Mathematics “BASIS”
1. Introduction
1.1.
The problem of completeness of translates of a single function in $L^{p}(\mathbb{R})$ spaces goes back to the classical Wiener’s theorems [Wie32], which characterize the functions whose translates span $L^{1}(\mathbb{R})$ or $L^{2}(\mathbb{R})$. If $g\in L^{1}(\mathbb{R})$ then the system of translates
$\{g(x-\lambda)\},\;\lambda\in\mathbb{R},$ (1.1)
is complete in the space $L^{1}(\mathbb{R})$ if and only if the Fourier transform $\widehat{g}$ has no zeros, while if $g\in L^{2}(\mathbb{R})$ then the system (1.1) is complete in $L^{2}(\mathbb{R})$ if and only if $\widehat{g}(t)\neq 0$ a.e.
Beurling [Beu51] proved that if $g\in(L^{p}\cap L^{1})(\mathbb{R})$, $21<p<$ 2, then the translates (1.1) span $L^{p}(\mathbb{R})$ if the zero set of $\widehat{g}$ has Hausdorff dimension less than $2(p-1)/p$. However, this sufficient condition is not necessary. Moreover, the functions $g$ whose translates span $L^{p}(\mathbb{R})$, $21<p<$ 2, cannot be characterized by the zero set of $\widehat{g}$, see [LO11].
1.2.
It is well known that even a discrete sequence of translates may suffice to span the space $L^{p}(\mathbb{R})$. Let us say that a discrete set $\Lambda\subset\mathbb{R}$ is $p$-generating if there exists a function $g\in L^{p}(\mathbb{R})$ (called a “generator”) such that the system of its $\Lambda$-translates
$\{g(x-\lambda)\},\;\lambda\in\Lambda,$ (1.2)
is complete in the space $L^{p}(\mathbb{R})$.
Which sets $\Lambda$ are $p$-generating? Note that if $\Lambda$ is $p$-generating for some $p$, then it is also $q$-generating for every $q>p$, see [OU16, Section 12.6]. It is thus “more difficult” to span the space $L^{p}(\mathbb{R})$ for smaller values of $p$.
For example, if $p>2$ then the set $\Lambda=\mathbb{Z}$ is $p$-generating, i.e. there is a function $g\in L^{p}(\mathbb{R})$ whose integer translates span $L^{p}(\mathbb{R})$. This fact was proved in [AO96]. The result was improved later: it follows from [FOSZ14, Theorem 3.2] that any unbounded set $\Lambda$, no matter how sparse, is $p$-generating for every $p>2$.
To the contrary, the set of integers $\Lambda=\mathbb{Z}$ is not $2$-generating. Indeed, by a simple argument involving the Fourier transform, one can show that a $2$-generating set $\Lambda$ cannot be contained in any arithmetic progression, see [OU16, Section 11.1].
However, it was proved in [Ole97] that any “almost integer” sequence
$\Lambda=\{n+\alpha_{n}:n\in\mathbb{Z}\},\quad 0\neq\alpha_{n}\to 0\quad(|n|\to+\infty),$ (1.3)
is $2$-generating. In particular, there exist uniformly discrete $2$-generating sets. (We recall that a set $\Lambda\subset\mathbb{R}$ is said to be uniformly discrete if the distance between any two distinct points of $\Lambda$ is bounded from below by a positive constant.)
The case $21<p<$ 2 seems to be more difficult, mainly due to the absence of Plancherel’s theorem in the corresponding $L^{p}(\mathbb{R})$ spaces. Only in [OU18], using methods of complex analysis, it was shown that if the perturbations $\alpha_{n}$ are exponentially small, then the set $\Lambda$ in (1.3) is $p$-generating for every $p>1$. A different approach, based on a result from [Lan64], was developed in [Lev25], which allows one to construct $p$-generating sets, $p>1$, consisting only of perturbations of the positive integers.
On the other hand, for $p=1$ there are no uniformly discrete generating sets. It was proved in [BOU06] that a set $\Lambda$ is $1$-generating if and only if the Beurling–Malliavin density of $\Lambda$ is infinite (for the definition of this density, see e.g. [OU16, Section 4.7]).
2. Results
2.1.
The results mentioned above provide a complete characterization of the discrete $p$-generating sets for $p=1$ and for $p>2$. To the contrary, the $p$-generating sets for $1<p\leqslant 2$, remain not well understood. In this case, both the size and the arithmetic structure of the set play an important role.
On one hand, a $p$-generating set, $1<p\leqslant 2$, cannot be too sparse. If a positive real sequence $\{\lambda_{n}\}_{n=1}^{\infty}$ satisfies the Hadamard lacunarity condition $\lambda_{n+1}/\lambda_{n}>c>1$, then it is not $2$-generating (and hence not $p$-generating for all $p\leqslant 2$), see [OU16, Section 11.4].
It is known [Ole98], [NO09] that this lacunarity condition is, in a sense, sharp: for any positive sequence $\varepsilon_{n}$ tending to 00, no matter how slowly, there exists a $2$-generating set $\Lambda=\{\lambda_{n}\}_{n=1}^{\infty}$ of positive real numbers satisfying $\lambda_{n+1}/\lambda_{n}>1+\varepsilon_{n}$ for all $n$. This result yields very sparse $2$-generating sets, having any subexponential growth.
Our first theorem extends the aforementioned result to the whole range of exponents $1<p\leqslant 2$, and moreover, we prove that the “generator” $g$ (namely, the function whose $\Lambda$-translates span the space) can be chosen to be a nonnegative function.
Theorem 2.1.
For any positive sequence $\varepsilon_{n}\to 0$ and any $\lambda_{0}>0$, one can find a nonnegative function $g\in\cap_{p>1}L^{p}(\mathbb{R})$ and a positive real sequence $\{\lambda_{n}\}_{n=1}^{\infty}$ satisfying
$\lambda_{n+1}/\lambda_{n}>1+\varepsilon_{n},\quad n=0,1,2,\dots,$ (2.1)
such that the system $\{g(x-\lambda_{n})\}_{n=1}^{\infty}$ is complete in the space $L^{p}(\mathbb{R})$ for every $p>1$.
2.2.
Next, we recall that not only the size, but also the arithmetic structure of the set is important. It was mentioned above that a $2$-generating set (and hence also a $p$-generating set, $1<p\leqslant 2$) cannot be contained in any arithmetic progression.
The question of whether every “almost integer” sequence of the form (1.3) is $p$-generating, $21<p<$ 2, was posed in [Ole97, Section 7] and has since remained open. Our second theorem provides the question with an affirmative answer and extends the result from [Ole97] to all values $1<p\leqslant 2$.
Theorem 2.2.
Let $\{\lambda_{n}\}_{n=1}^{\infty}$ be any real sequence satisfying $\lambda_{n}=n+\alpha_{n}$, where $0\neq\alpha_{n}\to 0$. Then there is a nonnegative function $g\in\cap_{p>1}L^{p}(\mathbb{R})$ such that the system $\{g(x-\lambda_{n})\}_{n=1}^{\infty}$ is complete in the space $L^{p}(\mathbb{R})$ for every $p>1$.
We note that in this result we use perturbations of the positive integers only.
2.3.
One can impose a more rigid structure on the set $\Lambda$ and require the successive differences $\lambda_{n+1}-\lambda_{n}$ to attain only finitely many different positive values. A real sequence $\{\lambda_{n}\}_{n=1}^{\infty}$ satisfying this condition is said to have finite local complexity.
Our third result shows that there exist $p$-generating sets of finite local complexity for every $p>1$, and moreover, the successive differences $\lambda_{n+1}-\lambda_{n}$ may attain as few as two different values.
Theorem 2.3.
Let $a,b>0$ be linearly independent over the rationals. Then there exist a real sequence $\{\lambda_{n}\}_{n=1}^{\infty}$ with $\lambda_{n+1}-\lambda_{n}\in\{a,b\}$, and a nonnegative Schwartz function $g$, such that the system $\{g(x-\lambda_{n})\}_{n=1}^{\infty}$ is complete in $L^{p}(\mathbb{R})$ for every $p>1$.
The requirement that $a,b$ be linearly independent over the rationals is crucial, for otherwise the points $\{\lambda_{n}\}$ lie in some arithmetic progression and the result fails.
Note that in Theorem 2.3 the function $g$ belongs to the Schwartz class, so it is both smooth and has fast decay. To the contrary, in Theorems 2.1 and 2.2 the function $g$ can be chosen smooth but in general cannot decay fast, see Sections 5.5 and 6.4.
2.4.
The rest of the paper is organized as follows. In Section 3 we recall some necessary background and fix notation that will be used throughout the paper.
In Section 4 we review and extend the approach from [Lev25], based on Landau’s classical result [Lan64], that allows one to construct uniformly discrete $p$-generating sets for every $p>1$ with a nonnegative Schwartz generator.
In Section 5 we construct a positive sequence $\{\lambda_{n}\}_{n=1}^{\infty}$ which is $p$-generating for every $p>1$, and the ratios $\lambda_{n+1}/\lambda_{n}$ tend to $1$ arbitrarily slowly (Theorem 2.1). In Section 6 we show that every “almost integer” sequence is $p$-generating for every $p>1$ (Theorem 2.2). Finally, in Section 7 we construct a sequence $\{\lambda_{n}\}_{n=1}^{\infty}$ which is $p$-generating for every $p>1$, and the differences $\lambda_{n+1}-\lambda_{n}$ attain only two different values (Theorem 2.3).
3. Preliminaries
In this section we recall some necessary background and fix notation that will be used throughout the paper.
3.1.
The Schwartz space $\mathcal{S}(\mathbb{R})$ consists of all infinitely smooth functions $\varphi$ on $\mathbb{R}$ such that for each $n,k\geqslant 0$, the seminorm
$\|\varphi\|_{n,k}:=\sup_{x\in\mathbb{R}}(1+|x|)^{n}|\varphi^{(k)}(x)|$ (3.1)
is finite. It is a topological linear space whose topology is induced by the metric
$d(\varphi,\psi):=\sum_{n,k\geqslant 0}2^{-(n+k)}\frac{\|\varphi-\psi\|_{n,k}}{1+\|\varphi-\psi\|_{n,k}}$ (3.2)
which also makes $\mathcal{S}(\mathbb{R})$ a complete, separable metric space.
A tempered distribution on $\mathbb{R}$ is a linear functional on the Schwartz space $\mathcal{S}(\mathbb{R})$ which is continuous with respect to the metric (3.2). We use $\alpha(\varphi)$ to denote the action of a tempered distribution $\alpha$ on a Schwartz function $\varphi$.
We denote by $\operatorname{supp}(\alpha)$ the closed support of a tempered distribution $\alpha$.
If $\varphi$ is a Schwartz function on $\mathbb{R}$ then we define its Fourier transform by
[…]
The Fourier transform of a tempered distribution $\alpha$ is defined by $\widehat{\alpha}(\varphi)=\alpha(\widehat{\varphi})$.
3.2.
Let $A^{p}(\mathbb{T})$, $1\leqslant p<\infty$, denote the Banach space of Schwartz distributions $\alpha$ on the circle $\mathbb{T}=\mathbb{R}/\mathbb{Z}$ whose Fourier coefficients $\{\widehat{\alpha}(n)\}$, $n\in\mathbb{Z}$, belong to $\ell^{p}(\mathbb{Z})$, endowed with the norm $\|\alpha\|_{A^{p}(\mathbb{T})}:=\|\widehat{\alpha}\|_{\ell^{p}(\mathbb{Z})}$. For $p=1$ this is the classical Wiener algebra $A(\mathbb{T})$ of continuous functions with an absolutely convergent Fourier series.
We also use $A^{p}(\mathbb{R})$, $1\leqslant p<\infty$, to denote the Banach space of tempered distributions $\alpha$ on $\mathbb{R}$ whose Fourier transform $\widehat{\alpha}$ is in $L^{p}(\mathbb{R})$, with the norm $\|\alpha\|_{A^{p}(\mathbb{R})}:=\|\widehat{\alpha}\|_{L^{p}(\mathbb{R})}$.
Note that $A^{1}(\mathbb{T})$ and $A^{1}(\mathbb{R})$ are function spaces, continuously embedded in $C(\mathbb{T})$ and $C_{0}(\mathbb{R})$ respectively. Similarly, for $1<p\leqslant 2$ the space $A^{p}$ (on either $\mathbb{T}$ or $\mathbb{R}$) is a function space, continuously embedded in $L^{q}$, $q=p/(p-1)$, by the Hausdorff-Young inequality. On the other hand, $A^{p}$ is not a function space for $p>2$.
3.3.
If $1\leqslant p<q<\infty$ then we have $A^{p}(\mathbb{T})\subset A^{q}(\mathbb{T})$, and moreover, the inequality
$\|\alpha\|_{A^{q}(\mathbb{T})}\leqslant\|\alpha\|_{A^{p}(\mathbb{T})}$ (3.4)
holds for every $\alpha\in A^{p}(\mathbb{T})$. To the contrary, there exists neither inclusion nor norm inequality between different $A^{p}(\mathbb{R})$ spaces.
If $\alpha\in A^{p}$ and $f\in A^{1}$ (on either $\mathbb{T}$ or $\mathbb{R}$) then the product $\alpha\cdot f$ is well defined, and
$\|\alpha\cdot f\|_{A^{p}}\leqslant\|\alpha\|_{A^{p}}\cdot\|f\|_{A^{1}}.$ (3.5)
3.4.
We introduce an auxiliary norm $\|{{\cdot}}\|_{{*}}$ on the Schwartz space $\mathcal{S}(\mathbb{R})$, defined by
$\|{{u}}\|_{{*}}:=10\cdot\sup_{x\in\mathbb{R}}(1+x^{2})|\widehat{u}(x)|,\quad u\in\mathcal{S}(\mathbb{R}).$ (3.6)
If $u\in\mathcal{S}(\mathbb{R})$, $f\in A^{p}(\mathbb{R})$ and $1\leqslant p\leqslant q<\infty$, then
$\|u\|_{A^{p}(\mathbb{R})}\leqslant\|{{u}}\|_{{*}},\quad\|u\cdot f\|_{A^{q}(\mathbb{R})}\leqslant\|{{u}}\|_{{*}}\|f\|_{A^{p}(\mathbb{R})}.$ (3.7)
Indeed, the first estimate is elementary while the second follows by an application of Young’s convolution inequality.
If $u\in\mathcal{S}(\mathbb{R})$ and $f\in A^{p}(\mathbb{T})$, $1\leqslant p<\infty$, then $f$ may be considered also as a $1$-periodic tempered distribution on $\mathbb{R}$, so the product $u\cdot f$ makes sense and is well defined.
Lemma 3.1.
Let $u\in\mathcal{S}(\mathbb{R})$ and $f\in A^{p}(\mathbb{T})$, $1\leqslant p<\infty$. Then
$\|u\cdot f\|_{A^{p}(\mathbb{R})}\leqslant\|{{u}}\|_{{*}}\|f\|_{A^{p}(\mathbb{T})}.$ (3.8)
For a proof see [LT24, Lemma 2.1].
3.5.
[…]
$\|\Delta_{h}\|_{A^{p}(\mathbb{T})}\leqslant h^{(p-1)/p},\quad 1\leqslant p<\infty.$ (3.9)
[…]
$\|\tau_{h}\|_{A^{p}(\mathbb{T})}\leqslant 3h^{(p-1)/p},\quad 1\leqslant p<\infty,$ (3.10)
which follows from (3.9) and the fact that $\tau_{h}(t)=\Delta_{h}(t+h)+\Delta_{h}(t)+\Delta_{h}(t-h)$.
3.6.
By a trigonometric polynomial we mean a finite sum of the form
$P(t)=\sum_{j}a_{j}e^{2\pi i\sigma_{j}t},\quad t\in\mathbb{R},$ (3.11)
where $\{\sigma_{j}\}$ are distinct real numbers, and $\{a_{j}\}$ are complex numbers.
By the spectrum of $P$ we mean the set $\operatorname{spec}(P):=\{\sigma_{j}:a_{j}\neq 0\}$. We observe that if $P$ has integer spectrum, $\operatorname{spec}(P)\subset\mathbb{Z}$, then $P$ is $1$-periodic, that is, $P(t+1)=P(t)$. In this case, $P$ may be considered also as a function on $\mathbb{T}$.
By the degree of $P$ we mean the number $\deg(P):=\min\{r\geqslant 0:\operatorname{spec}(P)\subset[-r,r]\}$.
For a trigonometric polynomial (3.11) we use the notation
$\|\widehat{P}\|_{p}=(\sum_{j}|a_{j}|^{p})^{1/p}.$ (3.12)
If $f\in A^{p}(\mathbb{R})$ and $P$ is a trigonometric polynomial, then
$\|f\cdot P\|_{A^{p}(\mathbb{R})}\leqslant\|\widehat{P}\|_{1}\cdot\|f\|_{A^{p}(\mathbb{R})}.$ (3.13)
4. Landau systems and completeness of weighted exponentials
In this section we review and extend the approach from [Lev25], based on Landau’s classical result [Lan64], which allows one to construct uniformly discrete $p$-generating sets $\Lambda$ for every $p>1$. Moreover, by some additional arguments we prove that the “generator” $g$ can be chosen to be a nonnegative Schwartz function.
4.1.
We begin by introducing some terminology.
Definition 4.1.
A set $\Omega\subset\mathbb{R}$ will be called a Landau set if it has the form
$\Omega=\Omega(L,h)=\bigcup_{|l|\leqslant L}[l-\tfrac{1}{2}h,l+\tfrac{1}{2}h]$ (4.1)
[…]
Thus a Landau set consists of a finite number of disjoint closed intervals of length strictly less than one, whose centers lie at integer points.
Definition 4.2.
A real sequence $\{\lambda_{n}\}_{n=1}^{\infty}$ will be called a Landau system if for every $N$ the system $\{e^{2\pi i\lambda_{n}t}\}$, $n>N$, is complete in $L^{2}(\Omega)$ for every Landau set $\Omega\subset\mathbb{R}$.
Landau [Lan64] proved that for any $\varepsilon>0$ there exists a Landau system $\{\lambda_{n}\}_{n=1}^{\infty}$ such that $|\lambda_{n}-n|<\varepsilon$ for all $n$. Using this result we can easily obtain:
Proposition 4.3.
There exists a Landau system $\{\lambda_{n}\}_{n=1}^{\infty}$ satisfying $\lambda_{n}=n+o(1)$.
Proof.
Indeed, let $\{\chi_{k}\}_{k=1}^{\infty}$ be a sequence dense in the Schwartz space, and fix a sequence of Landau sets $\Omega_{k}=\Omega(L_{k},h_{k})$ such that $L_{k}\to+\infty$ and $h_{k}\to 1$. By Landau’s theorem [Lan64], for each $k$ there is a Landau system $\{\lambda^{(k)}_{n}\}_{n=1}^{\infty}$ satisfying $|\lambda^{(k)}_{n}-n|<k^{-1}$ for all $n$. In particular, for every $N$ the system $\{e^{2\pi i\lambda^{(k)}_{n}t}\}$, $n>N$, is complete in $L^{2}(\Omega_{k})$. This allows us to construct by induction a sequence of trigonometric polynomials
$P_{k}(t)=\sum_{N_{k}<n\leqslant N_{k+1}}c_{n}e^{2\pi i\lambda^{(k)}_{n}t}$ (4.2)
such that $\|P_{k}-\chi_{k}\|_{L^{2}(\Omega_{k})}<k^{-1}$. In turn, we next construct a sequence $\{\lambda_{n}\}_{n=1}^{\infty}$ defined by $\lambda_{n}:=\lambda^{(k)}_{n}$ if $n$ is a positive integer belonging to the interval $(N_{k},N_{k+1}]$. It follows that $|\lambda_{n}-n|<k^{-1}$ for $n>N_{k}$, so that $\lambda_{n}=n+o(1)$.
Let us show that $\{\lambda_{n}\}_{n=1}^{\infty}$ is a Landau system. Indeed, let $\Omega=\Omega(L,h)$ be a Landau set. Then for all sufficiently large $k$ we have $\Omega\subset\Omega_{k}$, hence $\|P_{k}-\chi_{k}\|_{L^{2}(\Omega)}\leqslant k^{-1}$. The sequence $\{P_{k}\}$ is therefore dense in the space $L^{2}(\Omega)$. Moreover, for each $k$, the polynomial $P_{k}$ lies in the linear span of the system $\{e^{2\pi i\lambda_{n}t}\}$, $n>N_{k}$. This implies that for every $N$ the system $\{e^{2\pi i\lambda_{n}t}\}$, $n>N$, is complete in the space $L^{2}(\Omega)$. ∎
4.2.
We now come to the main result of the present section.
Definition 4.4.
We denote by $I_{0}(\mathbb{R})$ the closed linear subspace of the Schwartz space $\mathcal{S}(\mathbb{R})$, that consists of all the functions $\varphi\in\mathcal{S}(\mathbb{R})$ satisfying
$\varphi^{(j)}(n+\tfrac{1}{2})=0,\quad n\in\mathbb{Z},\quad j=0,1,2,\dots$ (4.3)
The space $I_{0}(\mathbb{R})$ is a complete, separable metric space with the metric inherited from the Schwartz space.
Theorem 4.5.
Given any Landau system $\{\lambda_{n}\}_{n=1}^{\infty}$ one can find a nonnegative function $u\in I_{0}(\mathbb{R})$, such that for every $N$ the system
$\{u(t)e^{2\pi i\lambda_{n}t}\},\quad n>N,$ (4.4)
is complete in $I_{0}(\mathbb{R})$. Moreover, $u$ can be chosen so that also $\widehat{u}$ is nonnegative.
This is a version of [Lev25, Theorem 4.2] where in addition both $u$ and $\widehat{u}$ are required to be nonnegative. Note that this extra nonnegativity requirement necessitates us to use here a slightly different definition of the space $I_{0}(\mathbb{R})$ than in [Lev25].
4.3.
Next, we turn to the proof of Theorem 4.5.
Let $J_{0}(\mathbb{R})$ be the linear space consisting of all the smooth, compactly supported functions on $\mathbb{R}$ which vanish in a neighborhood of $\mathbb{Z}+\tfrac{1}{2}$. Equivalently, $J_{0}(\mathbb{R})$ consists of all the smooth functions whose support is contained in some Landau set.
The space $J_{0}(\mathbb{R})$ is a dense subspace of $I_{0}(\mathbb{R})$, see [Lev25, Lemma 4.3].
Lemma 4.6.
Let $\chi\in J_{0}(\mathbb{R})$. Then there is a function $\sigma$ with the following properties:
  1. (i)
    $\sigma\in J_{0}(\mathbb{R})$;
  2. (ii)
    $\sigma(t)>0$ for $t\in\operatorname{supp}(\chi)$;
  3. (iii)
    $\sigma$ and $\widehat{\sigma}$ are both nonnegative functions.
Proof.
Let $\rho$ be an even smooth function on $\mathbb{R}$, with $\rho(t)>0$ for $|t|<\tfrac{1}{2}$, and $\rho(t)=0$ for $|t|\geqslant\tfrac{1}{2}$. By replacing $\rho(t)$ with $(\rho\ast\rho)(2t)$ we may assume that $\widehat{\rho}$ is nonnegative.
[…]
$\sigma(t):=\sum_{|l|\leqslant L}\Big{(}1-\frac{|l|}{L+1}\Big{)}\rho\Big{(}\frac{t-l}{h^{\prime}}\Big{)}.$ (4.5)
Then $\sigma$ is a smooth function supported on $\Omega(L,h^{\prime})$, hence (i) holds. Moreover, $\sigma(t)>0$ for $t\in\Omega(L,h)$, so that also condition (ii) is satisfied.
Lastly, it is obvious that $\sigma$ is a nonnegative function. Moreover,
$\widehat{\sigma}(x)=h^{\prime}\cdot\widehat{\rho}(h^{\prime}x)\sum_{|l|\leqslant L}\Big{(}1-\frac{|l|}{L+1}\Big{)}e^{2\pi ilx},$ (4.6)
and we observe that the sum in (4.6) is the classical $L$’th order Fejér kernel, which is nonnegative. Hence (iii) holds as well and the lemma is established. ∎
Proof of Theorem 4.5.
Let $\{\lambda_{n}\}_{n=1}^{\infty}$ be a Landau system, and choose a sequence $\{\chi_{k}\}_{k=1}^{\infty}$ in $J_{0}(\mathbb{R})$ which is dense in the space $I_{0}(\mathbb{R})$. We construct by induction a sequence of functions $u_{k}\in J_{0}(\mathbb{R})$, where $u_{k}$ and $\widehat{u}_{k}$ are both nonnegative, together with an increasing sequence of positive integers $\{N_{k}\}$, and trigonometric polynomials
$P_{k}(t)=\sum_{N_{k}<n<N_{k+1}}c_{n}e^{2\pi i\lambda_{n}t}$ (4.7)
in the following way. We begin by setting $u_{0}:=0$ and $N_{0}:=0$.
At the $k$’th step of the induction, we apply Lemma 4.6 with $\chi=\chi_{k}$ and obtain a function $\sigma_{k}$. Let $u_{k}:=u_{k-1}+\delta_{k}\sigma_{k}$, where $\delta_{k}>0$ is chosen small enough so that
$d(u_{k},u_{k-1})\leqslant 2^{-k},\quad\max_{1\leqslant l\leqslant k-1}d(u_{k}\cdot P_{l},u_{k-1}\cdot P_{l})\leqslant k^{-1}2^{-k}.$ (4.8)
We claim that there exists a polynomial $P_{k}$ as in (4.7), such that $d(u_{k}\cdot P_{k},\chi_{k})<k^{-1}$. Indeed, suppose that $\alpha$ is a tempered distribution on $\mathbb{R}$, which annihilates the system $\{u_{k}(t)e^{2\pi i\lambda_{n}t}\}$, $n>N_{k}$. This means that $\alpha\cdot u_{k}$ satisfies the condition $(\alpha\cdot u_{k})^{\wedge}(-\lambda_{n})=0$ for all $n>N_{k}$. Since $u_{k}\in J_{0}(\mathbb{R})$, the distribution $\alpha\cdot u_{k}$ is supported on some Landau set $\Omega_{k}$. Since $\{\lambda_{n}\}_{n=1}^{\infty}$ is a Landau system, this implies that $\alpha\cdot u_{k}=0$, see [Lev25, Corollary 3.2]. On the other hand, we note that $u_{k}(t)>0$ for $t\in\operatorname{supp}(\chi_{k})$, so we can write $\chi_{k}=u_{k}\cdot v_{k}$ where $v_{k}$ is a smooth function of compact support. Hence
$\alpha(\chi_{k})=\alpha(u_{k}\cdot v_{k})=(\alpha\cdot u_{k})(v_{k})=0.$ (4.9)
By duality, this implies that the function $\chi_{k}$ must lie in the closed linear subspace of $I_{0}(\mathbb{R})$ spanned by the system $\{u_{k}(t)e^{2\pi i\lambda_{n}t}\}$, $n>N_{k}$. As a consequence, we conclude that indeed there is a polynomial (4.7) satisfying $d(u_{k}\cdot P_{k},\chi_{k})<k^{-1}$.
The sequence $\{u_{k}\}$ converges in the space $I_{0}(\mathbb{R})$ to some function $u\in I_{0}(\mathbb{R})$ such that both $u$ and $\widehat{u}$ are nonnegative. We have
$d(u\cdot P_{k},\chi_{k})\leqslant d(u_{k}\cdot P_{k},\chi_{k})+\sum_{l=k+1}^{\infty}d(u_{l}\cdot P_{k},u_{l-1}\cdot P_{k})<2k^{-1}$ (4.10)
for every $k$, due to (4.8). As a consequence, the sequence $\{u\cdot P_{k}\}$ is dense in $I_{0}(\mathbb{R})$. Moreover, $u\cdot P_{k}$ belongs to the linear span of the system $\{u(t)e^{2\pi i\lambda_{n}t}\}$, $n>N_{k}$, due to (4.7). This implies that for every $N$ the system (4.4) is complete in $I_{0}(\mathbb{R})$. ∎
4.4.
As a consequence of the result just proved, we obtain:
Corollary 4.7.
Given any Landau system $\{\lambda_{n}\}_{n=1}^{\infty}$ one can find a nonnegative Schwartz function $g$ on $\mathbb{R}$, such that for every $N$ the system
$\{g(x-\lambda_{n})\},\quad n>N,$ (4.11)
is complete in the space $L^{p}(\mathbb{R})$ for every $p>1$.
Indeed, this follows from Theorem 4.5 and the fact that the Fourier transform maps $I_{0}(\mathbb{R})$ continuously and densely into $L^{p}(\mathbb{R})$, $p>1$, see [Lev25, Section 5].
In particular, by applying Corollary 4.7 to a Landau system $\{\lambda_{n}\}_{n=1}^{\infty}$ satisfying the condition $\lambda_{n}=n+o(1)$, we obtain [Lev25, Theorem 1.1] with the extra property that the “generator” $g$ is taken to be nonnegative.
5. Sparse complete sequences of translates
5.1.
In this section we prove Theorem 2.1. First, by using the fact that the Fourier transform is an isometric isomorphism $A^{p}(\mathbb{R})\to L^{p}(\mathbb{R})$, we can reformulate Theorem 2.1 as a result about completeness of weighted exponentials in $A^{p}(\mathbb{R})$.
Theorem 5.1.
For any positive sequence $\varepsilon_{n}\to 0$ and any $\lambda_{0}>0$, there is a nonnegative $w\in\cap_{p>1}A^{p}(\mathbb{R})$ with $\widehat{w}$ nonnegative, and a positive real sequence $\{\lambda_{n}\}_{n=1}^{\infty}$ satisfying
$\lambda_{n+1}/\lambda_{n}>1+\varepsilon_{n},\quad n=0,1,2,\dots,$ (5.1)
such that the system $\{w(t)e^{2\pi i\lambda_{n}t}\}_{n=1}^{\infty}$ is complete in the space $A^{p}(\mathbb{R})$ for every $p>1$.
Theorem 2.1 is obtained as a consequence of this result, by taking $g=\widehat{w}$.
The proof of Theorem 5.1 given below combines the approach in [NO09] together with techniques from [Lev25] and [LT24].
5.2.
Lemma 5.2.
Given any $60<h<1/$ 6, one can find a nonnegative function $\varphi\in A(\mathbb{T})$ with the following properties:
  1. (i)
    $\widehat{\varphi}(0)=1$, $\widehat{\varphi}(n)\geqslant 0$ for all $n\in\mathbb{Z}$;
  2. (ii)
[…]
  3. (iii)
    $\|\varphi-1\|_{A^{p}(\mathbb{T})}\leqslant 6\cdot h^{(p-1)/p}$ for every $p\geqslant 1$.
Proof.
It can be verified in a similar way to [LT24, Lemma 5.3] that the function
$\varphi(t):=1+3\cdot\Delta_{h}(t)-\tau_{h}(t-1/2)$ (5.2)
satisfies the required properties. ∎
Lemma 5.3.
Given $p>1$ and $\varepsilon>0$ there exist two trigonometric polynomials $P$ and $\gamma$ with integer spectrum, such that
  1. (i)
    $\gamma(t)>0$ for all $t\in\mathbb{T}$ (in particular, $\gamma$ has no zeros);
  2. (ii)
    $\widehat{\gamma}(0)=1$, $\widehat{\gamma}(n)\geqslant 0$ for all $n\in\mathbb{Z}$;
  3. (iii)
    $\|\gamma-1\|_{A^{p}(\mathbb{T})}<\varepsilon$;
  4. (iv)
    $\widehat{P}(-n)=0$ for $n=0,1,2,\dots$;
  5. (v)
    $\|P\cdot\gamma-1\|_{A^{p}(\mathbb{T})}<\varepsilon$.
Proof.
We choose and fix a small $h=h(p,\varepsilon)>0$. Let $\varphi$ be the function given by Lemma 5.2, and let $\chi(t):=1-\tau_{2h}(t-1/2)$. We first claim that one can find a trigonometric polynomial $P$ with integer spectrum satisfying the condition (iv) and such that $\|P\cdot\varphi-\chi\|_{A^{p}(\mathbb{T})}<\varepsilon/2$.
Indeed, let $\alpha$ be a Schwartz distribution belonging to the dual space $(A^{p}(\mathbb{T}))^{*}=A^{q}(\mathbb{T})$, $q=p/(p-1)$, and suppose that $\alpha$ annihilates the system
$\{\varphi(t)e^{2\pi int}\},\quad n=1,2,3,\dots$ (5.3)
[…]
$\alpha(\chi)=\alpha(\varphi\cdot\psi\cdot\chi)=(\alpha\cdot\varphi)(\psi\cdot\chi)=0.$ (5.4)
By duality, this implies that $\chi$ must lie in the closed linear subspace of $A^{p}(\mathbb{T})$ spanned by the system (5.3), so there is a trigonometric polynomial $P$ with integer spectrum satisfying (iv) and such that $\|P\cdot\varphi-\chi\|_{A^{p}(\mathbb{T})}<\varepsilon/2$.
If $h=h(p,\varepsilon)>0$ is small enough, then both $\|\varphi-1\|_{A^{p}(\mathbb{T})}$ and $\|\chi-1\|_{A^{p}(\mathbb{T})}$ do not exceed $\varepsilon/2$. As a consequence, $\|P\cdot\varphi-1\|_{A^{p}(\mathbb{T})}<\varepsilon$. We can thus conclude the proof by choosing $\gamma$ to be a Fejér sum of $\varphi$ of sufficiently high order, that is, $\gamma=\varphi\ast K_{N}$ where $K_{N}$ is the classical $N$’th order Fejér kernel and $N$ is sufficiently large. ∎
5.3.
Lemma 5.4.
Let $u\in\mathcal{S}(\mathbb{R})$, let $H$ be a trigonometric polynomial (with real spectrum), and let $p>1$ and $\eta>0$ be given. Then there is $\delta>0$ such that for any $d>0$ one can find trigonometric polynomials $\Gamma$ and $Q$ satisfying the following properties:
  1. (i)
    $\operatorname{spec}(\Gamma)\subset\mathbb{Z}$;
  2. (ii)
    $\Gamma(t)>0$ for all $t\in\mathbb{T}$ (in particular, $\Gamma$ has no zeros);
  3. (iii)
    $\widehat{\Gamma}(0)=1$, $\widehat{\Gamma}(n)\geqslant 0$ for all $n\in\mathbb{Z}$;
  4. (iv)
    $\|\Gamma-1\|_{A^{p}(\mathbb{T})}<\eta$;
  5. (v)
    $\operatorname{spec}(Q)$ is contained in some set $\{\lambda_{1},\lambda_{2},\dots,\lambda_{N}\}\subset\mathbb{R}$;
  6. (vi)
    $\lambda_{1}>d$, and $\lambda_{j+1}/\lambda_{j}>1+\delta$ for $j=1,2,\dots,N-1$;
  7. (vii)
    $\|u\cdot(\Gamma\cdot Q-H)\|_{A^{q}(\mathbb{R})}<\eta$ for every $q\geqslant p$.
Proof.
Let us denote $H(t)=\sum_{n=1}^{K}c_{n}e^{2\pi i\sigma_{n}t}$, where $\{\sigma_{n}\}$ are distinct real numbers, and $\{c_{n}\}$ are complex numbers. We choose a small $\varepsilon=\varepsilon(u,H,p,\eta)>0$, and apply Lemma 5.3 to obtain trigonometric polynomials $P$ and $\gamma$. Next, we choose large positive integers $\nu_{1}<\nu_{2}<\dots<\nu_{K}$ and set
$\Gamma(t)=\prod_{n=1}^{K}\gamma(\nu_{n}t),\quad Q(t)=\sum_{n=1}^{K}c_{n}e^{2\pi i\sigma_{n}t}P(\nu_{n}t).$ (5.5)
It is obvious that the properties (i) and (ii) are satisfied.
If $\nu_{1},\dots,\nu_{K}$ are chosen sufficiently fast increasing, then also (iii) holds. Moreover,
$\|\Gamma-1\|_{A^{p}(\mathbb{T})}^{p}=\|\Gamma\|_{A^{p}(\mathbb{T})}^{p}-1=\|\gamma\|_{A^{p}(\mathbb{T})}^{pK}-1<(1+\varepsilon)^{pK}-1<\eta^{p}$ (5.6)
provided that $\varepsilon$ is chosen small enough, so we obtain (iv).
(The second equality in (5.6) can be verified by expanding $\gamma$ in its (finite) Fourier series and opening the brackets; see [LT24, Lemma 3.2] which is similar.)
Let $M=M(u,H)=1+\sum_{n=1}^{K}\|{{c_{n}u(t)e^{2\pi i\sigma_{n}t}}}\|_{{*}}$. We claim that
$\|\Gamma(t)P(\nu_{n}t)-1\|_{A^{p}(\mathbb{T})}<M^{-1}\eta,\quad n=1,2,\dots,K,$ (5.7)
for $\varepsilon$ small enough. Indeed, we have
$\Gamma(t)P(\nu_{n}t)-1=\Gamma_{n}(t)(\gamma(\nu_{n}t)P(\nu_{n}t)-1)+(\Gamma_{n}(t)-1),$ (5.8)
where $\Gamma_{n}(t):=\prod_{j\neq n}\gamma(\nu_{j}t)$. The norm of the last summand in (5.8) can be estimated similarly to (5.6), so we can assume that $\|\Gamma_{n}-1\|_{A^{p}(\mathbb{T})}<(2M)^{-1}\eta$ by choosing $\varepsilon$ small enough. If $\nu_{1},\dots,\nu_{K}$ are chosen sufficiently fast increasing, then the norm of the first summand on the right hand side of (5.8) is equal to
$\|\Gamma_{n}\|_{A^{p}(\mathbb{T})}\cdot\|\gamma\cdot P-1\|_{A^{p}(\mathbb{T})}\leqslant 2\varepsilon<(2M)^{-1}\eta$ (5.9)
for $\varepsilon$ small enough, and thus (5.7) follows.
Next we verify property (vii). Indeed, we have
$u(t)(\Gamma(t)Q(t)-H(t))=\sum_{n=1}^{K}c_{n}u(t)e^{2\pi i\sigma_{n}t}(\Gamma(t)P(\nu_{n}t)-1),$ (5.10)
hence using (3.4), (3.8), (5.7) we obtain for every $q\geqslant p$,
$\|u\cdot(\Gamma\cdot Q-H)\|_{A^{q}(\mathbb{R})}\leqslant\sum_{n=1}^{K}\|{{c_{n}u(t)e^{2\pi i\sigma_{n}t}}}\|_{{*}}\cdot\|\Gamma(t)P(\nu_{n}t)-1\|_{A^{p}(\mathbb{T})}<\eta,$ (5.11)
so condition (vii) is satisfied.
Lastly we turn to establish the properties (v) and (vi). We denote $L=\deg(P)$, and take $\delta=(1+L)^{-1}$. Now suppose that $d>0$ is given. If $\nu_{1},\dots,\nu_{K}$ are chosen sufficiently fast increasing, then we have
$\operatorname{spec}(Q)\subset\bigcup_{n=1}^{K}J_{n},\quad J_{n}=\sigma_{n}+\{\nu_{n},2\nu_{n},3\nu_{n},\dots,L\nu_{n}\},$ (5.12)
and $J_{n+1}$ follows $J_{n}$ for each $n=1,2,\dots,K-1$. Let us write
$\bigcup_{n=1}^{K}J_{n}=\{\lambda_{1},\lambda_{2},\dots,\lambda_{N}\},$ (5.13)
where $\lambda_{j}$ are ordered increasingly. Then $\lambda_{1}=\sigma_{1}+\nu_{1}>d$, provided that $\nu_{1}$ is large enough (we note that $\nu_{1},\dots,\nu_{K}$ are allowed to depend on $d$, while $\delta$ does not depend on $\nu_{1},\dots,\nu_{K}$). Next, suppose that $\lambda_{j}$ and $\lambda_{j+1}$ are two consecutive elements of the set (5.13). If $\lambda_{j}$ and $\lambda_{j+1}$ lie in the same block $J_{n}$, then according to (5.12) we have
$\frac{\lambda_{j+1}}{\lambda_{j}}=\frac{\sigma_{n}+(l+1)\nu_{n}}{\sigma_{n}+l\nu_{n}}$ (5.14)
for some $l\in\{1,2,\dots,L-1\}$. Otherwise, $\lambda_{j}$ is the last element of $J_{n}$, while $\lambda_{j+1}$ is the first element of $J_{n+1}$, for some $n\in\{1,2,\dots,K-1\}$. In this case,
$\frac{\lambda_{j+1}}{\lambda_{j}}=\frac{\sigma_{n+1}+\nu_{n+1}}{\sigma_{n}+L\nu_{n}}.$ (5.15)
In either case, (5.14) or (5.15), we can ensure by taking $\nu_{1},\dots,\nu_{K}$ large enough that the condition $\lambda_{j+1}/\lambda_{j}>1+\delta$ holds for $j=1,2,\dots,N-1$. Thus properties (v) and (vi) are satisfied and the lemma is proved. ∎
5.4. Proof of Theorem 5.1
The proof consists of several steps.
5.4.1.
Let $0<\varepsilon_{n}\to 0$ be given, and we may also assume that $\varepsilon_{n+1}<\varepsilon_{n}$ for each $n$. We start by choosing a sequence $\{\chi_{k}\}_{k=1}^{\infty}\subset I_{0}(\mathbb{R})$ which is dense in the space $I_{0}(\mathbb{R})$.
By Proposition 4.3 there exists a Landau system $\{\sigma_{n}\}_{n=1}^{\infty}$ satisfying
$\sigma_{n}=n+o(1),\quad n\to+\infty.$ (5.16)
By an application of Theorem 4.5 we obtain a nonnegative function $u_{0}\in I_{0}(\mathbb{R})$ with $\widehat{u}_{0}$ nonnegative, such that for every $N$ the system $\{u_{0}(t)e^{2\pi i\sigma_{n}t}\}$, $n>N$, is complete in $I_{0}(\mathbb{R})$. We will construct by induction a sequence of functions $\{u_{k}\}_{k=1}^{\infty}\subset I_{0}(\mathbb{R})$ such that for each $k$ and every $N$, the system
$\{u_{k}(t)e^{2\pi i\sigma_{n}t}\},\;n>N,$ (5.17)
is complete in the space $I_{0}(\mathbb{R})$.
The sequence $\{\lambda_{n}\}_{n=1}^{\infty}$ satisfying (5.1) will be constructed by the same induction.
At the $k$’th step of the induction, suppose that the elements $\{\lambda_{n}\}$, $1\leqslant n\leqslant N_{k}$, satisfying (5.1) have already been defined. Given any positive integer $l_{k}$ we use the completeness of the system $\{u_{k-1}(t)e^{2\pi i\sigma_{n}t}\}$, $n>l_{k}$, in $I_{0}(\mathbb{R})$ to find a polynomial
$H_{k}(t)=\sum_{l_{k}<n<l^{\prime}_{k}}c_{n,k}e^{2\pi i\sigma_{n}t}$ (5.18)
such that
$d(u_{k-1}\cdot H_{k},\chi_{k})<k^{-1}.$ (5.19)
We choose a small number $\eta_{k}>0$ so that
$\eta_{k}\cdot\Big{(}1+\|{{u_{k-1}}}\|_{{*}}+\sum_{j=1}^{k-1}\|{{u_{k-1}\cdot Q_{j}}}\|_{{*}}\Big{)}<2^{-k}k^{-1},$ (5.20)
and let $p_{k}:=1+k^{-1}$. Now invoke Lemma 5.4 with $u_{k-1}$, $H_{k}$, $p_{k}$ and $\eta_{k}$ to obtain a number $\delta_{k}>0$. We choose $M_{k}$ large enough so that we have $\varepsilon_{M_{k}}<\delta_{k}$, and add arbitrary elements $\{\lambda_{n}\}$, $N_{k}<n\leqslant M_{k}$, while keeping the condition (5.1). Now set $d_{k}:=(1+\varepsilon_{M_{k}})\lambda_{M_{k}}$ and use Lemma 5.4 to obtain trigonometric polynomials $\Gamma_{k}$ and $Q_{k}$. The spectrum of $Q_{k}$ is contained in a sequence $\{\lambda_{n}\}$, $M_{k}<n\leqslant N_{k+1}$, which still satisfies (5.1) thanks to Lemma 5.4(v),(vi).
Finally, define $u_{k}:=u_{k-1}\cdot\Gamma_{k}$ which is a nonnegative function in $I_{0}(\mathbb{R})$. Recall that by the previous inductive step, for every $N$ the system $\{u_{k-1}(t)e^{2\pi i\sigma_{n}t}\}$, $n>N$, is complete in $I_{0}(\mathbb{R})$. Since multiplication by a trigonometric polynomial $\Gamma_{k}$ with no zeros maps $I_{0}(\mathbb{R})$ continuously and densely into $I_{0}(\mathbb{R})$, it follows that also the system (5.17) is complete in the space $I_{0}(\mathbb{R})$. This concludes the inductive construction.
5.4.2.
Since we have $1<p_{k}\to 1$, then for each $p>1$ there exists a sufficiently large $k(p)$ such that $1<p_{k}<p$ for all $k\geqslant k(p)$. For such $k$, using (3.4), (3.8) we have
$\|u_{k}-u_{k-1}\|_{A^{p}(\mathbb{R})}=\|u_{k-1}\cdot(\Gamma_{k}-1)\|_{A^{p}(\mathbb{R})}\leqslant\|{{u_{k-1}}}\|_{{*}}\|\Gamma_{k}-1\|_{A^{p_{k}}(\mathbb{T})}.$ (5.21)
Using Lemma 5.4(iv) and (5.20) we conclude that
$\|u_{k}-u_{k-1}\|_{A^{p}(\mathbb{R})}<2^{-k},\quad k\geqslant k(p),$ (5.22)
which implies that the sequence $u_{k}$ converges in the space $A^{p}(\mathbb{R})$ for every $p>1$.
The limit of the sequence $u_{k}$ is thus a nonnegative function $w\in\cap_{p>1}A^{p}(\mathbb{R})$. We recall that $u_{0}$ has a nonnegative Fourier transform $\widehat{u}_{0}$, while each $\Gamma_{k}$ is a trigonometric polynomial with nonnegative Fourier coefficients. This implies that $\widehat{u}_{k}$ is nonnegative for each $k$, and as a consequence, $\widehat{w}$ is a nonnegative function in $\cap_{p>1}L^{p}(\mathbb{R})$.
5.4.3.
Next, we fix $p>1$ and claim that for every $N$ the system
$\{w(t)e^{2\pi i\lambda_{n}t}\},\quad n>N,$ (5.23)
is complete in the space $A^{p}(\mathbb{R})$. To prove this, it will be enough to show that
$\|w\cdot Q_{k}-\chi_{k}\|_{A^{p}(\mathbb{R})}\to 0,\quad k\to+\infty.$ (5.24)
Indeed, the sequence $\{\chi_{k}\}_{k=1}^{\infty}$ is dense in the space $I_{0}(\mathbb{R})$, while $I_{0}(\mathbb{R})$ is continuously and densely embedded in $A^{p}(\mathbb{R})$, see [Lev25, Section 5]. Hence $\{\chi_{k}\}_{k=1}^{\infty}$ is dense also in the space $A^{p}(\mathbb{R})$. Recalling that $\operatorname{spec}(Q_{k})$ is contained in the sequence $\{\lambda_{n}\}$, $n>N_{k}$, the condition (5.24) implies the completeness of the system (5.23) in $A^{p}(\mathbb{R})$.
In order to establish (5.24) we write
$w\cdot Q_{k}-\chi_{k}=(u_{k-1}\cdot H_{k}-\chi_{k})+u_{k-1}\cdot(\Gamma_{k}\cdot Q_{k}-H_{k})+(w-u_{k})\cdot Q_{k},$ (5.25)
and estimate the $A^{p}(\mathbb{R})$ norm of each term on the right hand side of (5.25).
To estimate the norm of the first term, we recall that the space $I_{0}(\mathbb{R})$ is continuously embedded in $A^{p}(\mathbb{R})$, so (5.19) implies that
$\|u_{k-1}\cdot H_{k}-\chi_{k}\|_{A^{p}(\mathbb{R})}\to 0,\quad k\to+\infty.$ (5.26)
To estimate the norm of the second term we recall that by Lemma 5.4(vii) we have
$\|u_{k-1}\cdot(\Gamma_{k}\cdot Q_{k}-H_{k})\|_{A^{p}(\mathbb{R})}<\eta_{k},\quad k\geqslant k(p).$ (5.27)
It remains to estimate the last term in (5.25). We note that for any fixed $k$ we have $u_{j}\cdot Q_{k}\to w\cdot Q_{k}$ as $j\to+\infty$ in the $A^{p}(\mathbb{R})$ norm, hence for $k\geqslant k(p)$ we have
$\|(w-u_{k})\cdot Q_{k}\|_{A^{p}(\mathbb{R})}\leqslant\sum_{j=k+1}^{\infty}\|(u_{j}-u_{j-1})\cdot Q_{k}\|_{A^{p}(\mathbb{R})}$ (5.28)
$\qquad=\sum_{j=k+1}^{\infty}\|u_{j-1}\cdot(\Gamma_{j}-1)\cdot Q_{k}\|_{A^{p}(\mathbb{R})}$ (5.29)
$\qquad\leqslant\sum_{j=k+1}^{\infty}\|{{u_{j-1}\cdot Q_{k}}}\|_{{*}}\cdot\|\Gamma_{j}-1\|_{A^{p_{j}}(\mathbb{T})}$ (5.30)
$\qquad\leqslant\sum_{j=k+1}^{\infty}2^{-j}j^{-1}<k^{-1},$ (5.31)
where the inequality (5.30) follows from (3.4), (3.8), while to obtain the inequality (5.31) we have used Lemma 5.4(iv) and (5.20). Finally, combining (5.25), (5.26), (5.27) and (5.28)–(5.31) yields the required estimate (5.24), and so we establish the completeness of the system (5.23) in $A^{p}(\mathbb{R})$ for every $p>1$.
This completes the proof of Theorem 5.1, and as a consequence, Theorem 2.1 is also established. ∎
5.5. Remarks
[…]
$M_{n}:=\sup_{t\in\mathbb{R}}\sum_{j\in\mathbb{Z}}u_{0}(t-j)(1+|t-j|)^{n}$ (5.32)
for each $k$, and the desired conclusion follows by passing to the limit as $k\to+\infty$.
2. On the other hand, if $\{\varepsilon_{n}\}$ decrease slowly, then the function $g$ cannot have fast decay, in fact, $g$ cannot be chosen in $L^{1}(\mathbb{R})$. For otherwise, $\widehat{g}$ would be a continuous function and the exponential system $\{e^{2\pi i\lambda_{n}t}\}$ would be complete on some interval of positive length, which is not possible if $\{\lambda_{n}\}$ is sparse, see e.g. [OU16, Section 4.7].
3. It can be inferred from the proof of Lemma 5.4 that by choosing $l_{k}\to+\infty$ and using (5.16), (5.18), we can obtain a sequence $\{\lambda_{n}\}$ such that $\operatorname{dist}(\lambda_{n},\mathbb{Z})\to 0$.
6. Completeness of almost integer translates
6.1.
In this section we prove Theorem 2.2, i.e. for any real sequence $\{\lambda_{n}\}_{n=1}^{\infty}$ satisfying
$\lambda_{n}=n+\alpha_{n},\quad 0\neq\alpha_{n}\to 0,$ (6.1)
we construct a nonnegative function $g\in\cap_{p>1}L^{p}(\mathbb{R})$ such that the system of translates $\{g(x-\lambda_{n})\}_{n=1}^{\infty}$ is complete in $L^{p}(\mathbb{R})$ for every $p>1$.
As before, since the Fourier transform is an isometric isomorphism $A^{p}(\mathbb{R})\to L^{p}(\mathbb{R})$, we will obtain Theorem 2.2 as a consequence of the following result:
Theorem 6.1.
Given any real sequence $\{\lambda_{n}\}_{n=1}^{\infty}$ satisfying (6.1) there is a nonnegative function $w\in\cap_{p>1}A^{p}(\mathbb{R})$ with $\widehat{w}$ nonnegative, such that the system $\{w(t)e^{2\pi i\lambda_{n}t}\}_{n=1}^{\infty}$ is complete in the space $A^{p}(\mathbb{R})$ for every $p>1$.
The proof combines our techniques with the original ideas from [Ole97].
6.2.
We will obtain Theorem 6.1 as a consequence of the following lemma.
Lemma 6.2.
Let $\{\lambda_{n}\}_{n=1}^{\infty}$ be a real sequence satisfying (6.1), let $v\in J_{0}(\mathbb{R})$ and let $f$ be a smooth function on $\mathbb{R}$. Then for any $p>1$, $\varepsilon>0$ and any positive integer $N$, one can find trigonometric polynomials $\Gamma$ and $Q$ with the following properties:
  1. (i)
    $\operatorname{spec}(\Gamma)\subset\mathbb{Z}$;
  2. (ii)
    $\Gamma(t)>0$ for all $t\in\mathbb{T}$ (in particular, $\Gamma$ has no zeros);
  3. (iii)
    $\widehat{\Gamma}(0)=1$, $\widehat{\Gamma}(n)\geqslant 0$ for all $n\in\mathbb{Z}$;
  4. (iv)
    $\|\Gamma-1\|_{A^{p}(\mathbb{T})}<\varepsilon$;
  5. (v)
    $\operatorname{spec}(Q)$ is contained in the set $\{\lambda_{n}\}$, $n>N$;
  6. (vi)
    $\|v\cdot(\Gamma\cdot Q-f)\|_{A^{q}(\mathbb{R})}<\varepsilon$ for every $q\geqslant p$.
We recall that the space $J_{0}(\mathbb{R})$ was defined in Section 4.3.
First we explain how Lemma 6.2 implies Theorem 6.1. It is done using a procedure similar to the one used to prove Theorems 4.5 and 5.1, so we shall be more brief.
Proof of Theorem 6.1 using Lemma 6.2.
We fix a sequence $\{\chi_{k}\}_{k=1}^{\infty}\subset J_{0}(\mathbb{R})$ which is dense in the space $I_{0}(\mathbb{R})$, and let $p_{k}:=1+k^{-1}$. We construct by induction a sequence of nonnegative functions $\{u_{k}\}_{k=1}^{\infty}\subset J_{0}(\mathbb{R})$ with $\widehat{u}_{k}$ also nonnegative, and trigonometric polynomials
$Q_{k}(t)=\sum_{N_{k}<n<N_{k+1}}c_{n}e^{2\pi i\lambda_{n}t}$ (6.2)
such that
$\|u_{k}-u_{k-1}\|_{A^{p}(\mathbb{R})}<2^{-k},\quad p\geqslant p_{k},$ (6.3)
and
$\|u_{k}\cdot Q_{j}-\chi_{j}\|_{A^{p}(\mathbb{R})}<j^{-1},\quad p\geqslant p_{j},\quad j=1,2,\ldots,k.$ (6.4)
We begin by setting $u_{0}:=0$ and $N_{0}:=0$. Suppose now that we have already defined the functions $u_{1},\dots,u_{k-1}$ and the trigonometric polynomials $Q_{1},\dots,Q_{k-1}$ satisfying
$\|u_{k-1}\cdot Q_{j}-\chi_{j}\|_{A^{p}(\mathbb{R})}<j^{-1},\quad p\geqslant p_{j},\quad j=1,2,\ldots,k-1.$ (6.5)
We apply Lemma 4.6 to $\chi_{k}$ and obtain a function $\sigma_{k}\in J_{0}(\mathbb{R})$ such that $\sigma_{k}(t)>0$ for $t\in\operatorname{supp}(\chi_{k})$, and both $\sigma$ and $\widehat{\sigma}$ are nonnegative functions. Put $v_{k}=u_{k-1}+\delta_{k}\sigma_{k}$, where $\delta_{k}>0$ is chosen small enough so that
$\|v_{k}-u_{k-1}\|_{A^{p}(\mathbb{R})}<2^{-k-1},\quad p\geqslant 1,$ (6.6)
and the inequalities (6.5) still hold with the function $v_{k}$ instead of $u_{k-1}$, that is,
$\|v_{k}\cdot Q_{j}-\chi_{j}\|_{A^{p}(\mathbb{R})}<j^{-1},\quad p\geqslant p_{j},\quad j=1,2,\ldots,k-1.$ (6.7)
The estimates (6.6), (6.7) may be established using (3.7), (3.13). Now, since $v_{k}(t)>0$ for $t\in\operatorname{supp}(\chi_{k})$, we may write $\chi_{k}=v_{k}\cdot f_{k}$ where $f_{k}$ is a smooth function. We then invoke Lemma 6.2 with $v_{k}$, $f_{k}$, $p_{k}$, a small number $\varepsilon_{k}>0$ and $N_{k}$, and obtain trigonometric polynomials $\Gamma_{k}$ and $Q_{k}$. Define $u_{k}=v_{k}\cdot\Gamma_{k}$ and observe that $u_{k}\in J_{0}(\mathbb{R})$, and both $u_{k}$ and $\widehat{u}_{k}$ are nonnegative functions. Now, if $p\geqslant p_{k}$ then using (3.4), (3.8) we obtain
$\|u_{k}-v_{k}\|_{A^{p}(\mathbb{R})}=\|v_{k}\cdot(\Gamma_{k}-1)\|_{A^{p}(\mathbb{R})}\leqslant\|{{v_{k}}}\|_{{*}}\|\Gamma_{k}-1\|_{A^{p_{k}}(\mathbb{T})}\leqslant\varepsilon_{k}\|{{v_{k}}}\|_{{*}},$ (6.8)
so if $\varepsilon_{k}$ is small enough then
$\|u_{k}-v_{k}\|_{A^{p}(\mathbb{R})}<2^{-k-1},\quad p\geqslant p_{k}.$ (6.9)
[…]
$\|u_{k}\cdot Q_{k}-\chi_{k}\|_{A^{p}(\mathbb{R})}=\|v_{k}\cdot(\Gamma_{k}\cdot Q_{k}-f_{k})\|_{A^{p}(\mathbb{R})}<\varepsilon_{k},\quad p\geqslant p_{k},$ (6.10)
[…]
It now follows from (6.3) that the sequence $\{u_{k}\}$ converges in the space $A^{p}(\mathbb{R})$ for every $p>1$ to a nonnegative function $w\in\cap_{p>1}A^{p}(\mathbb{R})$ such that $\widehat{w}$ is also nonnegative. We may pass to the limit as $k\to+\infty$ in the estimate (6.4) and conclude that
$\|w\cdot Q_{j}-\chi_{j}\|_{A^{p}(\mathbb{R})}\leqslant j^{-1},\quad p\geqslant p_{j},\quad j=1,2,\dots.$ (6.11)
The sequence $\{\chi_{j}\}$ is dense in $I_{0}(\mathbb{R})$ and therefore in $A^{p}(\mathbb{R})$ for every $p>1$. In turn, due to (6.11) this implies that $\{w\cdot Q_{j}\}$ is again a dense sequence in $A^{p}(\mathbb{R})$. By (6.2) we conclude that for every $N$ the system $\{w(t)e^{2\pi i\lambda_{n}t}\}$, $n>N$, is complete in $A^{p}(\mathbb{R})$. ∎
We have thus shown that Theorem 6.1 is a consequence of Lemma 6.2, and in turn, Theorem 2.2 follows. It therefore remains to prove the lemma.
6.3. Proof of Lemma 6.2
The proof will be done in several steps.
6.3.1.
First we observe that with no loss of generality, we may assume that $\operatorname{supp}(v)$ is contained in $[0,+\infty)$. Indeed, if this is not the case then we may apply a translation to the right by a sufficiently large positive integer. Note that the function $v$ remains in $J_{0}(\mathbb{R})$ after an integer translation. Moreover, the properties (i)(iv) of the trigonometric polynomial $\Gamma$ remain invariant under integer translations, due to the periodicity of $\Gamma$, while the properties (v)(vi) remain invariant under arbitrary real translations.
We may therefore assume that $v$ is supported on a set of the form
$\Omega=\bigcup_{j=0}^{s-1}[j-\tfrac{1}{2}h,j+\tfrac{1}{2}h]$ (6.12)
[…]
6.3.2.
Let us choose $h^{\prime}$ such that $h<h^{\prime}<1$, and fix a smooth function $\Phi$ satisfying
[…]
We define
$\Theta(t)=\sum_{j=0}^{s-1}\Phi(t-j)$ (6.14)
and observe that $\Theta(t)=1$ on $\Omega$, while $\operatorname{supp}(\Theta)$ is contained in the set
$\Omega^{\prime}=\bigcup_{j=0}^{s-1}[j-\tfrac{1}{2}h^{\prime},j+\tfrac{1}{2}h^{\prime}].$ (6.15)
Next, we choose $h^{\prime\prime}$ such that $h<h^{\prime}<h^{\prime\prime}<1$, and fix a smooth function $\Psi$ satisfying
[…]
We note the following simple properties of the functions thus defined,
$\Psi\cdot\Phi=\Phi,\quad\Theta\cdot v=v,$ (6.17)
and
$\Psi(t)\Theta(t+j)=\Phi(t),\quad 0\leqslant j\leqslant s-1.$ (6.18)
6.3.3.
A key idea in [Ole97] involves iterations of the difference operator $\Delta$ defined by
$\Delta^{0}\phi=\phi,\quad(\Delta\phi)(t)=\phi(t+1)-\phi(t),\quad\Delta^{k}\phi=\Delta(\Delta^{k-1}\phi),$ (6.19)
where $\phi$ is an arbitrary function on $\mathbb{R}$.
Due to (6.1), for any trigonometric polynomial $q(t)=\sum_{n}a_{n}e^{2\pi i\lambda_{n}t}$ we have
$(\Delta^{k}q)(t)=\sum_{n}a_{n}(e^{2\pi i\alpha_{n}}-1)^{k}e^{2\pi i\lambda_{n}t}.$ (6.20)
In particular, $\Delta^{k}q$ is also a trigonometric polynomial, and $\operatorname{spec}(\Delta^{k}q)\subset\operatorname{spec}(q)$.
The following identity was stated in [Ole97],
$\phi(t+j)=\sum_{l=0}^{j}\binom{j}{l}(\Delta^{l}\phi)(t),\quad 0\leqslant j\leqslant s-1.$ (6.21)
For the reader’s benefit we include a quick proof of (6.21). Let $I$ denote the identity operator, $I\phi=\phi$, and set $(T\phi)(t)=\phi(t+1)$, then we have $T=I+\Delta$. Hence $T^{j}\phi=(I+\Delta)^{j}\phi$, and after opening the brackets one arrives at the identity (6.21).
6.3.4.
The next lemma is a consequence of (6.21) which allows us to estimate the $A^{p}(\mathbb{R})$ norm of a function $\phi$ in terms of its iterated differences.
Lemma 6.3.
Let $\phi$ be a smooth function supported in $\Omega^{\prime}$. Then
$\|\phi\|_{A^{p}(\mathbb{R})}\leqslant 2^{s}\max_{0\leqslant l\leqslant s-1}\|\Psi\cdot\Delta^{l}\phi\|_{A^{p}(\mathbb{R})}.$ (6.22)
Proof.
If we multiply both sides of (6.21) by $\Psi(t)$, take the $A^{p}(\mathbb{R})$ norm of both sides and apply the triangle inequality, then we obtain
$\|\Psi(t)\phi(t+j)\|_{A^{p}(\mathbb{R})}\leqslant 2^{j}\max_{0\leqslant l\leqslant j}\|\Psi\cdot\Delta^{l}\phi\|_{A^{p}(\mathbb{R})}.$ (6.23)
Since $\phi$ is supported in $\Omega^{\prime}$, we have $\phi(t)=\sum_{j=0}^{s-1}\Psi(t-j)\phi(t)$, hence
$\|\phi\|_{A^{p}(\mathbb{R})}\leqslant\sum_{j=0}^{s-1}\|\Psi(t-j)\phi(t)\|_{A^{p}(\mathbb{R})}=\sum_{j=0}^{s-1}\|\Psi(t)\phi(t+j)\|_{A^{p}(\mathbb{R})},$ (6.24)
where the last equality holds due to the invariance of the $A^{p}(\mathbb{R})$ norm under translation. Combining the estimates (6.23) and (6.24) implies (6.22). ∎
6.3.5.
The identity (6.21) admits the following inverse,
$(\Delta^{l}\phi)(t)=\sum_{j=0}^{l}\binom{l}{j}(-1)^{l-j}\phi(t+j),\quad 0\leqslant l\leqslant s-1,$ (6.25)
which can be established in a similar way to (6.21). It follows from (6.18), (6.25) that
$\Psi\cdot\Delta^{l}(\Theta\cdot\phi)=\Phi\cdot\Delta^{l}\phi,\quad 0\leqslant l\leqslant s-1,$ (6.26)
where $\phi$ is an arbitrary function on $\mathbb{R}$. The straightforward verification is left to the reader.
6.3.6.
Let us now explain the role of Lemma 6.3 in the proof. We will show that thanks to this lemma, the last condition (vi) of Lemma 6.2 can be obtained as a consequence of the following condition,
$\|\Phi\cdot\Delta^{l}(\Gamma\cdot Q-f)\|_{A^{p}(\mathbb{R})}<\delta,\quad 0\leqslant l\leqslant s-1,$ (6.27)
provided that $\delta=\delta(\varepsilon,v)>0$ is a sufficiently small number.
Indeed, assume that (6.27) holds and let us check that condition (vi) follows. Let $q\geqslant p$, then using (3.7), (6.17), (6.22), (6.26), (6.27), and since $\operatorname{supp}(\Theta)\subset\Omega^{\prime}$, we obtain
$\|v\cdot(\Gamma\cdot Q-f)\|_{A^{q}(\mathbb{R})}=\|\Theta\cdot v\cdot(\Gamma\cdot Q-f)\|_{A^{q}(\mathbb{R})}$ (6.28)
$\qquad\leqslant\|{{v}}\|_{{*}}\cdot\|\Theta\cdot(\Gamma\cdot Q-f)\|_{A^{p}(\mathbb{R})}$ (6.29)
$\qquad\leqslant\|{{v}}\|_{{*}}\cdot 2^{s}\max_{0\leqslant l\leqslant s-1}\|\Psi\cdot\Delta^{l}(\Theta\cdot(\Gamma\cdot Q-f))\|_{A^{p}(\mathbb{R})}$ (6.30)
$\qquad=\|{{v}}\|_{{*}}\cdot 2^{s}\max_{0\leqslant l\leqslant s-1}\|\Phi\cdot\Delta^{l}(\Gamma\cdot Q-f)\|_{A^{p}(\mathbb{R})}\leqslant 2^{s}\delta\|{{v}}\|_{{*}},$ (6.31)
hence (vi) follows provided that $2^{s}\delta\|{{v}}\|_{{*}}<\varepsilon$.
6.3.7.
We now turn to the construction of the trigonometric polynomials $\Gamma$ and $Q$ satisfying the conditions (i)(vi) of Lemma 6.2. As we have just seen, the last condition (vi) will follow from (6.27).
The construction involves an inductive process that we now describe. By induction we will define trigonometric polynomials $q_{1},q_{2},\ldots,q_{s}$ such that
$q_{j}(t)=\sum_{N_{j}<n<N^{\prime}_{j}}c_{n}e^{2\pi i\lambda_{n}t},$ (6.32)
together with trigonometric polynomials $\gamma_{1},\gamma_{2},\ldots,\gamma_{s}$ satisfying
$\gamma_{j}(t)>0,\quad\operatorname{spec}({\gamma}_{j})\subset\mathbb{Z},\quad\widehat{\gamma}_{j}(0)=1,\quad\widehat{\gamma}_{j}(n)\geqslant 0,\quad n\in\mathbb{Z},$ (6.33)
and fast increasing integers $M_{1}<M_{2}<\dots<M_{s}$. We denote
$\gamma_{j}^{(M_{j})}(t)=\gamma_{j}(M_{j}t),\quad\Gamma_{l}(t)=\prod_{j=1}^{l}\gamma_{j}^{(M_{j})}(t),\quad Q_{l}(t)=\sum_{j=1}^{l}q_{j}(t).$ (6.34)
We also set $\Gamma_{0}=1$ and $Q_{0}=0$.
Suppose that at the $l$’th step we are given a small $\delta_{l}>0$ (the choice of $\delta_{l}$ will be specified later). We will construct $q_{l}$, $\gamma_{l}$ and $M_{l}$ so that the following properties hold,
$\|\Phi\cdot\Delta^{s-l}(f-\Gamma_{l}\cdot Q_{l})\|_{A^{p}(\mathbb{R})}<\delta_{l},$ (6.35)
$\|\Phi\cdot\Delta^{s-j}(\Gamma_{l}\cdot Q_{l}-\Gamma_{l-1}\cdot Q_{l-1})\|_{A^{p}(\mathbb{R})}<\delta_{l},\quad 1\leqslant j\leqslant l-1,$ (6.36)
$\|\Gamma_{l}-\Gamma_{l-1}\|_{A^{p}(\mathbb{T})}<\delta_{l}.$ (6.37)
The construction of $q_{l}$, $\gamma_{l}$ and $M_{l}$ is done as follows. Suppose that $q_{j}$, $\gamma_{j}$ and $M_{j}$ are already defined for $j=1,\dots,l-1$. Define a $1$-periodic function $g_{l}$ on $\mathbb{R}$ by
$g_{l}(t)=\frac{\Psi(t)}{\Gamma_{l-1}(t)}\cdot(\Delta^{s-l}(f-\Gamma_{l-1}\cdot Q_{l-1}))(t),\quad t\in[-\tfrac{1}{2},\tfrac{1}{2}).$ (6.38)
The fact that $\Psi$ is supported on $[-\tfrac{1}{2}h^{\prime\prime},\tfrac{1}{2}h^{\prime\prime}]$ implies that $g_{l}$ is a smooth function. In particular $g_{l}$ has an absolutely convergent Fourier series, that is, $g_{l}\in A(\mathbb{T})$. Let $\widetilde{g}_{l}$ be a Fourier partial sum of $g_{l}$ of sufficiently high order so that
$\|{{\Phi\cdot\Gamma_{l-1}}}\|_{{*}}\cdot\|g_{l}-\widetilde{g}_{l}\|_{A(\mathbb{T})}<\tfrac{1}{6}\delta_{l}.$ (6.39)
Now we choose a small number $\varepsilon_{l}>0$ and apply Lemma 5.3 with $\varepsilon=\varepsilon_{l}$. The lemma provides us with trigonometric polynomials $P_{l}$ and $\gamma_{l}$ with integer spectrum, such that $\gamma_{l}(t)>0$, $\widehat{\gamma}_{l}(0)=1$, $\widehat{\gamma}_{l}(n)\geqslant 0$ for $n\in\mathbb{Z}$, $\widehat{P}_{l}(-n)=0$ for $n\geqslant 0$, and
$\|\gamma_{l}-1\|_{A^{p}(\mathbb{T})}<\varepsilon_{l},\quad\|P_{l}\cdot\gamma_{l}-1\|_{A^{p}(\mathbb{T})}<\varepsilon_{l}.$ (6.40)
Next we choose a large positive integer $M_{l}$ and define a trigonometric polynomial
$p_{l}(t)=\widetilde{g}_{l}(t)\cdot P_{l}(M_{l}t).$ (6.41)
For any given integer $N_{l}$, we can choose $M_{l}=M_{l}(\widetilde{g}_{l},P_{l},N_{l})$ sufficiently large so that $\operatorname{spec}(p_{l})$ contains only frequencies larger than $N_{l}$, so that $p_{l}$ has the form
$p_{l}(t)=\sum_{N_{l}<n<N^{\prime}_{l}}b_{n}e^{2\pi int}.$ (6.42)
We have the estimate
$\sum_{N_{l}<n<N^{\prime}_{l}}|b_{n}|=\|p_{l}\|_{A(\mathbb{T})}\leqslant\|g_{l}\|_{A(\mathbb{T})}\cdot\|P_{l}\|_{A(\mathbb{T})},$ (6.43)
and we note that this estimate does not depend on the choice of $N_{l}$ and $M_{l}$.
Now we define the trigonometric polynomial $q_{l}$ as follows,
$q_{l}(t)=\sum_{N_{l}<n<N^{\prime}_{l}}c_{n}e^{2\pi i\lambda_{n}t},\quad c_{n}=\frac{b_{n}}{(e^{2\pi i\alpha_{n}}-1)^{s-l}}.$ (6.44)
By choosing $N_{l}$ large enough and using the assumption $0\neq\alpha_{n}\to 0$, we can ensure that the denominator in (6.44) does not vanish. It follows from (6.20) that
$(\Delta^{s-j}q_{l})(t)=\sum_{N_{l}<n<N^{\prime}_{l}}b_{n}(e^{2\pi i\alpha_{n}}-1)^{l-j}e^{2\pi i\lambda_{n}t},\quad 1\leqslant j\leqslant l.$ (6.45)
In particular, for $j=l$ this yields
$(\Delta^{s-l}q_{l})(t)=\sum_{N_{l}<n<N^{\prime}_{l}}b_{n}e^{2\pi i\lambda_{n}t}.$ (6.46)
On the other hand, for $11\leqslant j\leqslant l-$ 1 we obtain from (6.43), (6.45) the following estimate for the $\ell^{1}$ norm of the coefficients of the trigonometric polynomial $\Delta^{s-j}q_{l}$,
$\|(\Delta^{s-j}q_{l})^{\wedge}\|_{1}\leqslant\|g_{l}\|_{A(\mathbb{T})}\cdot\|P_{l}\|_{A(\mathbb{T})}\cdot 2\pi\sup_{n>N_{l}}|\alpha_{n}|^{l-j},$ (6.47)
where we have also used here the elementary inequality $|e^{2\pi i\alpha_{n}}-1|\leqslant 2\pi|\alpha_{n}|$. Since $\alpha_{n}\to 0$, the right-hand side of (6.47) can be made arbitrarily small if $N_{l}$ is chosen large enough. Hence we may choose $N_{l}$ so that
$\|(\Delta^{s-j}q_{l})^{\wedge}\|_{1}<\varepsilon_{l},\quad 1\leqslant j\leqslant l-1.$ (6.48)
Let us now show that we can choose $\varepsilon_{l}$ small enough and $N_{l}$, $M_{l}$ large enough so that the conditions (6.35), (6.36), (6.37) are fulfilled. First we establish (6.37) as follows,
$\|\Gamma_{l}-\Gamma_{l-1}\|_{A^{p}(\mathbb{T})}=\|\Gamma_{l-1}\cdot(\gamma_{l}^{(M_{l})}-1)\|_{A^{p}(\mathbb{T})}$ (6.49)
$\qquad\leqslant\|\Gamma_{l-1}\|_{A(\mathbb{T})}\cdot\|\gamma_{l}-1\|_{A^{p}(\mathbb{T})}\leqslant\|\Gamma_{l-1}\|_{A(\mathbb{T})}\cdot\varepsilon_{l}<\delta_{l}$ (6.50)
provided that $\varepsilon_{l}$ is small enough.
Next we check the condition (6.36). We have
$\Gamma_{l}\cdot Q_{l}-\Gamma_{l-1}\cdot Q_{l-1}=\Gamma_{l-1}\cdot Q_{l-1}\cdot(\gamma_{l}^{(M_{l})}-1)+\Gamma_{l}\cdot q_{l}.$ (6.51)
We consider each of the two summands on the right hand side of (6.51). We start with the first summand. Note that if $\phi,\psi$ are two functions on $\mathbb{R}$ and $\phi$ is $1$-periodic, then $\Delta^{k}(\phi\cdot\psi)=\phi\cdot\Delta^{k}\psi$, which can be verified e.g. using (6.25). Hence
$\Phi\cdot\Delta^{s-j}(\Gamma_{l-1}\cdot Q_{l-1}\cdot(\gamma_{l}^{(M_{l})}-1))=\Phi\cdot\Gamma_{l-1}\cdot(\gamma_{l}^{(M_{l})}-1)\cdot\Delta^{s-j}(Q_{l-1}),$ (6.52)
and as a consequence, the $A^{p}(\mathbb{R})$ norm of (6.52) does not exceed
$\|{{\Phi\cdot\Gamma_{l-1}\cdot\Delta^{s-j}(Q_{l-1})}}\|_{{*}}\cdot\|\gamma_{l}-1\|_{A^{p}(\mathbb{T})}\leqslant\|{{\Phi\cdot\Gamma_{l-1}\cdot\Delta^{s-j}(Q_{l-1})}}\|_{{*}}\cdot\varepsilon_{l}<\tfrac{1}{2}\delta_{l}$ (6.53)
for $\varepsilon_{l}$ small enough. We next consider the second summand in (6.51). Since we have
$\Phi\cdot\Delta^{s-j}(\Gamma_{l}\cdot q_{l})=\Phi\cdot\Gamma_{l-1}\cdot\gamma_{l}^{(M_{l})}\cdot\Delta^{s-j}(q_{l}),$ (6.54)
and using (3.13), (6.48) we obtain that the $A^{p}(\mathbb{R})$ norm of (6.54) does not exceed
$\|{{\Phi\cdot\Gamma_{l-1}}}\|_{{*}}\cdot\|\gamma_{l}\|_{A^{p}(\mathbb{T})}\cdot\|(\Delta^{s-j}q_{l})^{\wedge}\|_{1}\leqslant\|{{\Phi\cdot\Gamma_{l-1}}}\|_{{*}}\cdot(1+\varepsilon_{l})\cdot\varepsilon_{l}<\tfrac{1}{2}\delta_{l}$ (6.55)
for $11\leqslant j\leqslant l-$ 1, provided that $\varepsilon_{l}$ is small enough. Hence we obtain (6.36) as a consequence of (6.51), (6.52), (6.54) and the estimates (6.53), (6.55).
Lastly, we must verify condition (6.35) as well. Due to (6.17), (6.38) we have
$\Phi\cdot\Gamma_{l-1}\cdot g_{l}=\Phi\cdot\Delta^{s-l}(f-\Gamma_{l-1}\cdot Q_{l-1}),$ (6.56)
hence
$\Phi\cdot\Delta^{s-l}(f-\Gamma_{l}\cdot Q_{l})=\Phi\cdot\Gamma_{l-1}\cdot g_{l}+\Phi\cdot\Delta^{s-l}(\Gamma_{l-1}\cdot Q_{l-1}-\Gamma_{l}\cdot Q_{l})$ (6.57)
$\qquad=\Phi\cdot\Gamma_{l-1}\cdot(g_{l}+\Delta^{s-l}(Q_{l-1}-\gamma_{l}^{(M_{l})}\cdot Q_{l}))$ (6.58)
$\qquad=\Phi\cdot\Gamma_{l-1}\cdot(g_{l}-\gamma_{l}^{(M_{l})}\cdot\Delta^{s-l}(q_{l}))+\Phi\cdot\Gamma_{l-1}\cdot(1-\gamma_{l}^{(M_{l})})\cdot\Delta^{s-l}(Q_{l-1}).$ (6.59)
We estimate the $A^{p}(\mathbb{R})$ norm of each summand in (6.59). The $A^{p}(\mathbb{R})$ norm of the second summand does not exceed
$\|{{\Phi\cdot\Gamma_{l-1}\cdot\Delta^{s-l}(Q_{l-1})}}\|_{{*}}\cdot\|1-\gamma_{l}\|_{A^{p}(\mathbb{T})}\leqslant\|{{\Phi\cdot\Gamma_{l-1}\cdot\Delta^{s-l}(Q_{l-1})}}\|_{{*}}\cdot\varepsilon_{l}<\tfrac{1}{2}\delta_{l}$ (6.60)
for $\varepsilon_{l}$ small enough.
To estimate the first summand in (6.59) we denote $P_{l}^{(M_{l})}(t)=P_{l}(M_{l}t)$, and recall that we have $p_{l}=\widetilde{g}_{l}\cdot P_{l}^{(M_{l})}$ according to (6.41). Hence we may write
$g_{l}-\gamma_{l}^{(M_{l})}\cdot\Delta^{s-l}(q_{l})=(g_{l}-\widetilde{g}_{l})+\widetilde{g}_{l}\cdot(1-P_{l}^{(M_{l})}\cdot\gamma_{l}^{(M_{l})})$ (6.61)
$\qquad+\gamma_{l}^{(M_{l})}\cdot(p_{l}-\Delta^{s-l}(q_{l})),$ (6.62)
and so now we have three terms to estimate. Using (3.4), (3.8) and (6.39) we have
$\|\Phi\cdot\Gamma_{l-1}\cdot(g_{l}-\widetilde{g}_{l})\|_{A^{p}(\mathbb{R})}\leqslant\|{{\Phi\cdot\Gamma_{l-1}}}\|_{{*}}\cdot\|g_{l}-\widetilde{g}_{l}\|_{A(\mathbb{T})}<\tfrac{1}{6}\delta_{l}.$ (6.63)
Next, we have
$\|\Phi\cdot\Gamma_{l-1}\cdot\widetilde{g}_{l}\cdot(1-P_{l}^{(M_{l})}\cdot\gamma_{l}^{(M_{l})})\|_{A^{p}(\mathbb{R})}$ (6.64)
$\qquad\leqslant\|{{\Phi\cdot\Gamma_{l-1}\cdot\widetilde{g}_{l}}}\|_{{*}}\cdot\|1-P_{l}\cdot\gamma_{l}\|_{A^{p}(\mathbb{T})}\leqslant\|{{\Phi\cdot\Gamma_{l-1}\cdot\widetilde{g}_{l}}}\|_{{*}}\cdot\varepsilon_{l}<\tfrac{1}{6}\delta_{l}$ (6.65)
for $\varepsilon_{l}$ small enough. Lastly, using (3.13), (6.42), (6.46), we have
$\|\Phi\cdot\Gamma_{l-1}\cdot\gamma_{l}^{(M_{l})}\cdot(p_{l}-\Delta^{s-l}(q_{l}))\|_{A^{p}(\mathbb{R})}$ (6.66)
$\qquad=\|\Phi(t)\cdot\Gamma_{l}(t)\cdot\sum_{N_{l}<n<N^{\prime}_{l}}b_{n}(e^{2\pi int}-e^{2\pi i\lambda_{n}t})\|_{A^{p}(\mathbb{R})}$ (6.67)
[…]
Note that the first two factors in (6.68) can be estimated independently of $N_{l}$ and $M_{l}$. Indeed, the first factor does not exceed $\|\Gamma_{l-1}\|_{A(\mathbb{T})}\cdot\|\gamma_{l}\|_{A(\mathbb{T})}$, while the second factor admits the estimate (6.43). Since we have $\lambda_{n}=n+\alpha_{n}$ and $\alpha_{n}\to 0$, we may therefore choose $N_{l}$ large enough so that (6.68) becomes smaller than $\frac{1}{6}\delta_{l}$. Summing up the last three estimates, and using (6.61)–(6.62), we obtain
$\|\Phi\cdot\Gamma_{l-1}\cdot(g_{l}-\gamma_{l}^{(M_{l})}\cdot\Delta^{s-l}(q_{l}))\|_{A^{p}(\mathbb{R})}<\tfrac{1}{2}\delta_{l}.$ (6.69)
In turn, together with (6.57)–(6.60) this finally yields the condition (6.35). Hence we have verified that our inductive procedure indeed provides us with $q_{l}$, $\gamma_{l}$ and $M_{l}$ such that (6.32), (6.33) hold and the properties (6.35), (6.36), (6.37) are satisfied.
6.3.8.
We are now able to complete the proof of Lemma 6.2 by exhibiting the required trigonometric polynomials $\Gamma$ and $Q$. Indeed, we take $\Gamma=\Gamma_{s}$ and $Q=Q_{s}$. We show that if at the $l$’th step of the inductive construction we choose $\delta_{l}$ sufficiently small and $N_{l}$, $M_{l}$ sufficiently large, then $\Gamma$ and $Q$ will satisfy conditions (i)(vi) of Lemma 6.2.
The fact that conditions (i)(ii) hold is obvious.
Since $\Gamma(t)=\prod_{l=1}^{s}\gamma_{l}(M_{l}t)$, if we choose $M_{l}$ at the $l$’th step large enough, then
$\widehat{\Gamma}(0)=\prod_{l=1}^{s}\widehat{\gamma}_{l}(0)=1.$ (6.70)
This can be verified, as before, by expanding each $\gamma_{l}$ in its (finite) Fourier series and opening the brackets, see [LT24, Lemma 3.2].
It is also obvious that $\widehat{\Gamma}(n)\geqslant 0$ for all $n\in\mathbb{Z}$, so we obtain condition (iii).
Next we note, using (6.37), that
$\|\Gamma-1\|_{A^{p}(\mathbb{T})}\leqslant\sum_{l=1}^{s}\|\Gamma_{l}-\Gamma_{l-1}\|_{A^{p}(\mathbb{T})}\leqslant\sum_{l=1}^{s}\delta_{l}<\varepsilon$ (6.71)
if the numbers $\delta_{l}$ are small enough, so the condition (iv) is established.
Since we have $Q(t)=\sum_{l=1}^{s}q_{l}(t)$, the condition (v) follows from (6.44) provided that all the numbers $N_{l}$ are larger than $N$.
It remains to verify the last condition (vi). We have seen above that this condition follows if (6.27) holds for a sufficiently small $\delta=\delta(\varepsilon,v)>0$. So, it suffices to show that
$\|\Phi\cdot\Delta^{s-l}(\Gamma\cdot Q-f)\|_{A^{p}(\mathbb{R})}<\delta,\quad 1\leqslant l\leqslant s,$ (6.72)
provided that the numbers $\delta_{l}$ are small enough. Indeed, due to (6.35), (6.36) we have
$\|\Phi\cdot\Delta^{s-l}(\Gamma\cdot Q-f)\|_{A^{p}(\mathbb{R})}\leqslant\|\Phi\cdot\Delta^{s-l}(\Gamma_{l}\cdot Q_{l}-f)\|_{A^{p}(\mathbb{R})}$ (6.73)
$\qquad+\sum_{k=l+1}^{s}\|\Phi\cdot\Delta^{s-l}(\Gamma_{k}\cdot Q_{k}-\Gamma_{k-1}\cdot Q_{k-1})\|_{A^{p}(\mathbb{R})}\leqslant\sum_{k=l}^{s}\delta_{l}<\delta,$ (6.74)
for $\delta_{l}$ small enough. Hence we verified that (6.27) holds, and as a consequence, the condition (vi) is established and Lemma 6.2 is proved. ∎
6.4. Remarks
[…]
$\sum_{j=1}^{k}\delta_{j}M_{n}(\sigma_{j}),\quad M_{n}(\sigma_{j}):=\sup_{t\in\mathbb{R}}\sum_{\nu\in\mathbb{Z}}\sigma_{j}(t-\nu)(1+|t-\nu|)^{n},$ (6.75)
so the desired conclusion follows e.g. by choosing $\delta_{j}$ such that $\delta_{j}M_{j}(\sigma_{j})<2^{-j}$.
2. The question whether $g$ may be chosen to have fast decay, was studied in [OU04]. It was proved that in general the answer is negative: there exists a real sequence $\{\lambda_{n}\}$, $n\in\mathbb{Z}$, such that $\lambda_{n}=n+\alpha_{n}$ where $0\neq\alpha_{n}\to 0$ as $|n|\to+\infty$, but such that there is no function $g\in(L^{2}\cap L^{1})(\mathbb{R})$ whose translates $\{g(x-\lambda_{n})\}$, $n\in\mathbb{Z}$, span $L^{2}(\mathbb{R})$. It follows, see [OU16, Proposition 12.24], that the same “almost integer” sequence $\{\lambda_{n}\}$ does not admit a generator $g\in(L^{p}\cap L^{1})(\mathbb{R})$ for the space $L^{p}(\mathbb{R})$, $21<p<$ 2.
7. Finite local complexity sequences of translates
7.1.
In this section we prove Theorem 2.3. The main ingredients of the proof consist of Theorem 4.5 (or its consequence Corollary 4.7), and the following lemma.
Lemma 7.1.
Let $a$ be an irrational positive real number, $\Omega=\Omega(L,h)$ be a Landau set given by (4.1), $\chi$ be a continuous function on $\Omega$, and let $\lambda_{0}\in\mathbb{R}$ and $\varepsilon>0$ be given. Then there exists a trigonometric polynomial $P(t)=\sum_{n=1}^{K}c_{n}e^{2\pi i\lambda_{n}t}$ such that
  1. (i)
    $|P(t)-\chi(t)|\leqslant\varepsilon$ for all $t\in\Omega$;
  2. (ii)
    $\lambda_{n+1}-\lambda_{n}\in\{1,a\}$ for $n=0,1,\dots,K-1$.
Proof.
By multiplying $\chi(t)$ on the exponential $e^{-2\pi i\lambda_{0}t}$ we may assume that $\lambda_{0}=0$. We will construct a trigonometric polynomial $P$ satisfying (i) which is of the form
$P(t)=\sum_{k=0}^{2L}e^{2\pi ikat}Q_{k}(t),$ (7.1)
where $Q_{k}$ are trigonometric polynomials with integer spectrum such that
$\operatorname{spec}(Q_{k})\subset\{j:N_{k}\leqslant j\leqslant N_{k+1}\},\quad 0\leqslant k\leqslant 2L,$ (7.2)
and $\{N_{k}\}$, $10\leqslant k\leqslant 2L+$ 1, is an increasing sequence of integers with $N_{0}=1$. Then
$\operatorname{spec}(P)\subset\bigcup_{k=0}^{2L}J_{k},\quad J_{k}=\{ka+j:\;N_{k}\leqslant j\leqslant N_{k+1}\},$ (7.3)
and $J_{k+1}$ follows $J_{k}$ for each $k=0,1,\dots,2L-1$. Let us write
$\bigcup_{k=0}^{2L}J_{k}=\{\lambda_{1},\lambda_{2},\dots,\lambda_{K}\},$ (7.4)
where $\lambda_{n}$ are ordered increasingly. Then $\lambda_{1}-\lambda_{0}=1$. Next, suppose that $\lambda_{n}$ and $\lambda_{n+1}$ are two consecutive elements of the set (7.4). If $\lambda_{n}$ and $\lambda_{n+1}$ lie in the same block $J_{k}$, then $\lambda_{n+1}-\lambda_{n}=1$. Otherwise, $\lambda_{n}$ is the last element of $J_{k}$, while $\lambda_{n+1}$ is the first element of $J_{k+1}$, for some $k\in\{0,1,\dots,2L-1\}$, and in this case $\lambda_{n+1}-\lambda_{n}=a$. We thus see that condition (ii) is satisfied.
We now turn to the construction of the polynomial $P$. First, we note that (7.1) and the periodicity of $Q_{k}$ imply that for any integer $l$ we have
$P(t+l)=\sum_{k=0}^{2L}e^{2\pi ikla}H_{k}(t),\quad H_{k}(t):=e^{2\pi ikat}Q_{k}(t).$ (7.5)
If we write (7.5) for $|l|\leqslant L$, then we obtain a system of $12L+$ 1 linear equations for the polynomials $H_{k}$, $0\leqslant k\leqslant 2L$, with a Vandermonde coefficient matrix $\{e^{2\pi ikla}\}$ whose determinant does not vanish due to the irrationality of $a$. By solving this system we get
$H_{k}(t)=\sum_{|l|\leqslant L}d_{kl}P(t+l),\quad 0\leqslant k\leqslant 2L,$ (7.6)
where $\{d_{kl}\}$ are the entries of the inverse matrix, which depend on $a$ and $L$ only.
Now recall that $10<h<$ 1, and hence for every $N$ the exponential system $\{e^{2\pi int}\}$, $n>N$, is complete in the space $C[-\tfrac{1}{2}h,\tfrac{1}{2}h]$, see e.g. [You01, Section 3.1]. Using this fact, we can choose the trigonometric polynomials $Q_{0},Q_{1},\ldots,Q_{2L}$ one after another, so that their spectra “follow one another”, i.e.  satisfy (7.2), and such that
$\max_{|t|\leqslant\frac{1}{2}h}\Big{|}Q_{k}(t)-e^{-2\pi ikat}\sum_{|l|\leqslant L}d_{kl}\chi(t+l)\Big{|}\leqslant\delta,\quad 0\leqslant k\leqslant 2L,$ (7.7)
where $\delta:=\varepsilon\cdot(2L+1)^{-1}$. Having chosen the polynomials $Q_{k}$, we next define the polynomial $P$ using (7.1). It now follows from (7.5), (7.6), (7.7) that if we denote
$E_{k}(t):=\sum_{|l|\leqslant L}d_{kl}(P(t+l)-\chi(t+l)),\quad t\in[-\tfrac{1}{2}h,\tfrac{1}{2}h],\quad 0\leqslant k\leqslant 2L,$ (7.8)
then $|E_{k}(t)|\leqslant\delta$. In turn, this implies that for $|l|\leqslant L$ and $t\in[-\tfrac{1}{2}h,\tfrac{1}{2}h]$ we have
$|P(t+l)-\chi(t+l)|=\Big{|}\sum_{k=0}^{2L}e^{2\pi ikla}E_{k}(t)\Big{|}\leqslant(2L+1)\cdot\delta=\varepsilon,$ (7.9)
that is, $|P(t)-\chi(t)|\leqslant\varepsilon$ for all $t\in\Omega$. Thus the lemma is proved. ∎
7.2. Proof of Theorem 2.3
By rescaling, it would be enough to prove the result in the case where one of the numbers $a,b$ is equal to $1$. Hence, in what follows we shall assume that $b=1$, and $a$ is an irrational positive number (since $a,b$ are linearly independent over the rationals).
By virtue of Corollary 4.7, it would suffice that we prove the existence of a Landau system $\{\lambda_{n}\}_{n=1}^{\infty}$ satisfying $\lambda_{n+1}-\lambda_{n}\in\{1,a\}$ for all $n$. Let $\{\chi_{k}\}_{k=1}^{\infty}$ be a sequence which is dense in the Schwartz space, and fix a sequence of Landau sets $\Omega_{k}=\Omega(L_{k},h_{k})$ such that $L_{k}\to+\infty$ and $h_{k}\to 1$. We now construct by induction an increasing sequence of positive integers $\{N_{k}\}$, and trigonometric polynomials
$P_{k}(t)=\sum_{N_{k}<n\leqslant N_{k+1}}c_{n}e^{2\pi i\lambda_{n}t}$ (7.10)
in the following way. At the $k$’th step of the induction, we apply Lemma 7.1 with $\Omega_{k}$, $\chi_{k}$ and $\varepsilon_{k}=k^{-1}$, in order to obtain a trigonometric polynomial $P_{k}$ of the form (7.10) such that $|P_{k}(t)-\chi_{k}(t)|\leqslant k^{-1}$ for all $t\in\Omega_{k}$. Moreover, Lemma 7.1 allows us to choose the frequencies $\{\lambda_{n}\}_{n=1}^{\infty}$ so that $\lambda_{n+1}-\lambda_{n}\in\{1,a\}$ for all $n$.
It remains to show that $\{\lambda_{n}\}_{n=1}^{\infty}$ is a Landau system. This can be done in the same way as in the proof of Proposition 4.3. Indeed, let $\Omega=\Omega(L,h)$ be a Landau set. Then for all sufficiently large $k$ we have $\Omega\subset\Omega_{k}$, hence $|P_{k}(t)-\chi_{k}(t)|\leqslant k^{-1}$ for all $t\in\Omega$. The sequence $\{P_{k}\}$ is therefore dense in the space $L^{2}(\Omega)$. Moreover, for each $k$, the polynomial $P_{k}$ lies in the linear span of the system $\{e^{2\pi i\lambda_{n}t}\}$, $n>N_{k}$. This implies that for every $N$ the system $\{e^{2\pi i\lambda_{n}t}\}$, $n>N$, is complete in the space $L^{2}(\Omega)$. Hence the sequence $\{\lambda_{n}\}_{n=1}^{\infty}$ is a Landau system, and Theorem 2.3 is thus proved. ∎